### v1

# Zeal: Rethinking Large-Scale Resource Allocation with “Decouple and Decompose”

Abstract
Resource allocation is fundamental for cloud systems to ensure efficient resource sharing among tenants. However, the scale of such optimization problems has outgrown the capabilities of commercial solvers traditionally employed in production. To scale up resource allocation, prior approaches either tailor solutions to specific problems or rely on assumptions tied to particular workloads. In this work, we revisit real-world resource allocation problems and uncover a common underlying structure: a vast majority of these problems are inherently separable, i.e., they optimize the aggregate utility of individual resource and demand allocations, under separate constraints for each resource and each demand. Building on this insight, we develop Zeal, a general, scalable, and theoretically grounded framework for accelerating resource allocation through a “decouple and decompose” approach. Zeal systematically decouples entangled resource and demand constraints, thereby decomposing the overall optimization into alternating per-resource and per-demand allocations, which can then be solved efficiently and in parallel. We have implemented Zeal as a library extension to an open-source solver, maintaining a familiar user interface. Experimental results across three prominent resource allocation tasks—traffic engineering, cluster scheduling, and load balancing—demonstrate Zeal’s substantial speedups and robust allocation quality.
1 Introduction
Resource allocation is critical to the orchestration of multi-tenant systems, which are managed by today’s cloud providers on an unprecedented scale. It allows cloud services to be delivered with efficiency, reliability, and cost-effectiveness, by judiciously distributing various compute resources (e.g., CPUs and GPUs) and I/O resources (e.g., memory, storage, and network bandwidth) across a vast number of clients.
Common applications of resource allocation include cluster scheduling [46, 27, 44, 61, 64], traffic engineering [24, 1, 67, 50, 31], and load balancing [16, 48, 52, 37]. To meet varying user or application demands (jobs, traffic flows, and workloads), these systems extensively employ commercial optimization solvers like Gurobi [22] to compute the allocation of expensive cloud infrastructure (cluster resources, network links, and servers) worth millions or even billions of dollars.
As cloud environments rapidly expand and diversify, the sheer scale of modern resource allocation problems has exceeded the capabilities of commercial solvers. These problems may involve millions of variables, which require solvers tens of minutes to hours to compute a solution when fast allocation is needed to maintain service quality. This “scalability crisis” in resource allocation has prompted a flurry of recent research [1, 43, 63, 47, 42] that trades off allocation quality for speed using heuristics, approximate algorithms, or machine learning. Nevertheless, these studies are limited to either a single domain (e.g., wide-area network traffic engineering [1, 63]) or a specific objective (e.g., max-min fairness [42]), requiring substantial manual effort to adapt to new scenarios.
In contrast, POP [43] takes a general approach to accelerating assorted resource allocation problems. It partitions a large problem into a few (e.g., 16) subproblems, each with a subset of resources and demands, and solves them independently. However, this partition hinges on a “granular” assumption—each demand requests only a small fraction of mutually interchangeable resources. This assumption proves brittle with real workloads and degrades solution quality (§7.2).
In this study, we re-examine real-world resource allocation problems through a different lens and aim to develop a general and scalable optimization framework without catering to specific problems or making workload-dependent assumptions. Our approach, Zeal, is theoretically grounded and enables fine-grained decomposition of the optimization problem, alternating between per-resource and per-demand allocations that can be solved in parallel (Figure 1).
Figure 1: Overview of Zeal.
This decomposition is made possible by an intrinsic, workload-agnostic problem structure that we uncovered after surveying dozens of real-world resource allocation problems in recent literature (Table 1). We discovered that a vast majority of these problems are inherently separable: they optimize the sum (or maximum) of utilities (measurement of solution quality) from allocating each individual resource and demand, while ensuring feasible allocation for individual resource and demand as well.111 We focus on individually separable resources and demands for simplicity and due to their widespread presence in real systems, but they can be relaxed to groups of resources and demands as we show in §4.1. In other words, the optimization objective comprises separate per-resource and/or per-demand utilities, and the constraints are also separate—they are either (1) a “resource constraint,” enforcing that the total demands mapped to each resource do not exceed the resource’s capacity, or (2) a “demand constraint,” enforcing that the total resources assigned to each demand do not surpass the demand’s needs. A formal definition of this separable structure is provided in §2, where we also show that real-world resource allocation is largely formulated with amenable mathematical properties (e.g., convex utilities, linear constraints) for tractability. [^2] Rare exceptions to this common structure are discussed in §4.2.
Nonetheless, the uncovered “separability” in resource allocation problems does not directly allow us to actually separate them into independent subproblems that address individual resources and demands as depicted in Figure 1. This is because each resource not only appears in its own resource constraint but also participates in the demand constraints associated with every demand mapped to it; likewise for each demand. Ignoring the dependencies across subproblems and greedily solving each subproblem independently would fail to find feasible and high-quality allocations.
Given the intertwined nature of resource and demand constraints, Zeal first decouples them as follows. Let $x$ denote the resource allocation matrix between $n$ resources and $m$ demands, where $x_{ij}$ represents the amount of resource $i$ assigned to demand $j$. The key mathematical trick in Zeal is to introduce a duplicate of the matrix $x$, denoted as $z$, along with an additional constraint $x=z$. Next, Zeal iteratively and alternately solves for $x$ subject only to resource constraints with $z$ fixed, and solves for $z$ subject only to demand constraints with $x$ fixed. In each iteration, Zeal further ensures that $x$ and $z$ do not deviate significantly from each other in order to satisfy $x=z$. This decoupling transformation lends itself to a classical constrained optimization framework called ADMM (Alternating Direction Method of Multipliers) [9]. ADMM allows Zeal to embed each of the resource and demand constraints into the original objective as a penalty for constraint violation, thereby constructing augmented Lagrangians [23] as new optimization objectives. ADMM provides the theoretical foundation that enables alternating optimizations of these augmented Lagrangians with respect to $x$ and $z$, respectively.
Next, Zeal further decomposes the above optimizations into per-resource and per-demand allocations. This is feasible owing to the separable structure we identified across resource allocation problems: the original objective is separable on each resource and demand variable, and so are the penalty terms converted by ADMM from constraints in the augmented Lagrangians. After decomposition, per-resource allocations involve solving $n$ subproblems, each with $m$ demand variables, and $m$ subproblems with $n$ resource variables for per-demand allocations. Consequently, the scale of optimization is greatly reduced from $n\times m$ variables to only $n$ or $m$ variables in each subproblem. These subproblems can be solved independently with massive parallelism using off-the-shelf solvers. Figure 2 illustrates the “decouple and decompose” workflow of Zeal.
To demonstrate Zeal’s practicality, we have implemented Zeal as a Python package [^3] Zeal can be installed with pip install zeal after publication., built on the widely used open-source solver cvxpy [17, 4]. The interface of Zeal is also designed to resemble cvxpy for ease of use (§6). Moreover, instead of simulating the parallel execution of subproblems on multiple CPU cores as previous studies did [1, 43], we have carefully optimized Zeal’s implementation using Ray [40], enabling efficient utilization of multiple cores without incurring significant overheads, such as cache contention.
We evaluate Zeal in three prominent resource allocation problems—traffic engineering, cluster scheduling, and load balancing—and demonstrate its faster speed and higher allocation quality than the state of the art (§7). Compared with the best variant of POP [43] for each problem, Zeal achieves a 5.3% higher allocation quality and a 7.6$\times$ speedup in traffic engineering, 14% and 1.34$\times$ in cluster scheduling, and 5.3% and 1.13$\times$ in load balancing.
Figure 2: Workflow of Zeal.
2 Real-World Resource Allocation Problems
In modern multi-tenant systems, resources such as CPUs, memory, and bandwidth are shared among numerous users and applications. A central resource allocator often casts the distribution of resources as an optimization problem (e.g., linear programming) and repeatedly computes feasible solutions with off-the-shelf solvers to accommodate changing demands.
However, these solvers are struggling to keep up with the increasing problem sizes that frequently occur in fast-growing systems. They may take several hours to solve an allocation problem that attempts to assign thousands of resources to thousands of demands—several orders of magnitude slower than desired (e.g., second-level SLAs, or service-level agreements). This challenge of scaling up resource allocation has spurred several recent works [1, 43, 63, 47, 42], but they either rely on domain-specific customizations or workload-dependent assumptions. In this paper, we instead seek to take a general (i.e., domain- and workload-agnostic) approach to scaling resource allocation.
After surveying real-world resource allocation problems from recent literature (Table 1), we discover that nearly all these formulations can be transformed into a separable structure (§1), which we formally describe as follows.
  • Variables.  An allocation matrix $x\in\mathcal{X}$ specifies how $n$ resources should be shared among $m$ demands, where $x_{ij}$ represents the quantity (or fraction) of resource $i$ assigned to demand $j$. Accordingly, the $i$th row of $x$, denoted as $x_{i*}=[x_{i1},x_{i2},\dots,x_{im}]$, constitutes the allocation vector of resource $i$, and the $j$th column $x_{*j}=[x_{1j},x_{2j},\dots,x_{nj}]$ encodes the allocation of demand $j$. The constraint set $\mathcal{X}$ is the Cartesian product of each $\mathcal{X}_{ij}$, which can be a set of real numbers or integers.
  • Objective.  The objective specifies a metric to be optimized over an allocation matrix $x$, primarily consisting of a sum of utilities (or costs, to measure allocation quality) from allocating each resource ($x_{i*}$) and demand ($x_{*j}$):
    $\sum_{i}f_{i}(x_{i*})+\sum_{j}g_{j}(x_{*j}),$ (1)
    where $f_{i}(\cdot):\mathbb{R}^{m}\rightarrow\mathbb{R}$ and $g_{j}(\cdot):\mathbb{R}^{n}\rightarrow\mathbb{R}$ are utility (or cost) functions. They are typically designed to be convex to ensure practical tractability. Minimizing the maximum of utilities can also be converted into regular minimizing utilities (Equation 1) with additional constraints [5].
  • Constraints.  To ensure that individual resources and demands are not over-provisioned (i.e., exceeding a resource’s capacity or a demand’s requirements), there is typically a constraint for each resource and each demand. We observe that these constraints are largely linear, presumably for tractability as well. For simplicity, we further convert inequality constraints into equality constraints using non-negative slack variables (details in Section 6). As such, the resource constraint on resource $i$ can be expressed as
    $R_{i}x_{i*}=r_{i},$ (2)
    and the demand constraint on demand $j$ can be written as
    $D_{j}x_{*j}=d_{j}.$ (3)
Despite the separable objective and separable constraints over each resource and each demand, large resource allocation problems (e.g., with millions of variables) still remain painfully slow to solve as they cannot actually be “separated” into smaller, concurrent per-resource and per-demand subproblems. The main obstacle is that resource and demand constraints are fundamentally coupled. Equations 2 and 3 show that each $x_{i,j}$ not only participates in the resource constraint (over $x_{i*}$) but also in the demand constraint (over $x_{*j}$). Next, we describe our solution to this challenge.
3 Zeal: Decouple and Decompose
We present Zeal, a general, scalable, and theoretically rooted framework for optimizing real-world resource allocation problems. Leveraging the separable structure (§2) that underpins a vast majority of these problems, Zeal enables massive parallelism with near-optimal solutions through a technique that we call “decouple and decompose.” A formal, mathematical representation of Zeal is included in Appendix A.
3.1 Decouple
Resource and demand constraints being intertwined is the fundamental cause of complexity in resource allocation, as the individual allocation of any resource or demand will affect all other allocations. An optimal solution requires collectively allocating all resources to all demands, while a greedy strategy will lead to suboptimal or unfeasible solutions.
Existing domain-agnostic solutions do not address this intrinsic problem nature of resource allocation. Production systems employ commercial solvers and rely on linear programming (LP) or mixed-integer linear programming (MILP) algorithms (e.g., simplex methods [45], barrier methods [30], or branch-and-bound methods [8]) to iteratively explore the feasible region—a high-dimensional search space dictated by coupled resource and demand constraints.
The state of the art, POP [43], naively breaks intertwined constraints by partitioning resources and demands into a handful of subsets (e.g., 4–64), and then independently allocating each subset of resources to the demands within the same subset. Doing so ignores the dependencies across subsets and leads to suboptimal solutions in practice (§7.1). Meanwhile, it only enables limited parallelism because finer-grained decomposition would render the allocation unfeasible and unusable.
Instead of disregarding the constraints across subproblems, we tackle the intertwining root cause directly. Specifically, we aim to decouple demand constraints from resource constraints, transforming the full allocation into alternating per-resource allocations—assigning a resource to all demands subject to the corresponding resource constraint, and per-demand allocations—mapping a demand to all resources are subject to the corresponding demand constraint.
To begin with, we offload demand constraints to another variable different from $x$. We create a copy of the allocation matrix $x$, denoted as $z$, and replace the variables in each demand constraint (Equation 3) with their counterparts in $z$:
$D_{j}z_{*j}=d_{j}.$ (4)
We also substitute $z$ for the demand-related utilities in the objective function (Equation 1):
$\sum_{i}f_{i}(x_{i*})+\sum_{j}g_{j}(z_{*j}).$ (5)
Then, representing the same allocation plan, $x$ and $z$ should be equal. Thus, we introduce a new constraint:
$x=z.$ (6)
Consequently, all resource constraints remain on $x$, but all demand constraints have been moved to $z$.
The decoupling of constraints clearly leads to a formulation equivalent to the original optimization with the new constraint $x=z$. However, the new constraint still has $x$ and $z$ entangled, so the introduction of $z$ does not immediately enable decomposition. In fact, jointly optimizing $x$ and $z$ (e.g., using penalty methods [6] or the augmented Lagrangian method [23]) will forfeit the benefits of decoupling, as we will show in §7.3.
On the bright side, the transformed formulation becomes compatible with a half-century-old constrained optimization framework called ADMM (Alternating Direction Method of Multipliers) [9], which provides theoretically guaranteed solutions when solving for two sets (“blocks”) of variables ($x$ and $z$). At a high level, the theory behind ADMM enables alternating optimization of the following subproblems without making compromises on the solution quality: minimizing the updated objective (Equation 5) with respect to
  • $x$ (with $z$ fixed), subject to resource constraints and $x=z$,
  • $z$ (with $x$ fixed), subject to demand constraints and $x=z$.
We elaborate in §3.2 that both optimization subproblems are amenable to decomposition, reducing problem sizes and enabling massive parallelism. Next, we first describe how ADMM solves them (even without decomposition).
  • Constraint transformation.  ADMM is a variant of the augmented Lagrangian method [23]. Taking the first subproblem above (for $x$) as an example: ADMM converts every constraint, i.e., resource constraints and $x=z$, into penalty terms and appends them to the objective to minimize (Equation 5). A conventional penalty method will add a single penalty term $\sum_{i}\mu\cdot c_{i}^{2}$ to the objective, where $c_{i}$ is a measure of the constraint violation (comprising $|R_{i}x_{i*}-r_{i}|$ and $|x_{i}-z_{i}|$) for resource $i$, and $\mu$ is a penalty coefficient that increases to infinity over iterations to ensure asymptotic convergence to the optimal solution. However, taking $\mu$ to infinity requires careful manual tuning, or otherwise incurs ill-conditioning [2]. The augmented Lagrangian method has a similar rationale to penalty methods, except that it introduces yet another penalty term $\lambda_{i}\cdot c_{i}$ to mimic a Lagrange multiplier. The update of $\lambda_{i}$ is automated as $\lambda_{i}\leftarrow\lambda_{i}+\mu\cdot c_{i}$ with theoretical guarantees, where $\mu$ can stay as a small constant while avoiding ill-conditioning.
  • Partial optimization.  The new objectives with additional penalty terms converted from constraint violations are also called augmented Lagrangians. ADMM alternately minimizes these augmented Lagrangians, solving for $x$ with $z$ fixed, and solving for $z$ with $x$ fixed. This is known as “two-block partial optimization” and guarantees the convergence to optimal solutions for convex problems [21, 20]. For non-convex problems (including those with integer variables), practical evidence suggests that ADMM is still effective at solving (some of) them as well [53, 62, 68]. Theoretical proofs are available under specific assumptions [56, 38, 60].
3.2 Decompose
To this point, the original resource allocation problem has been transformed into alternating minimization of two augmented Lagrangians (original objective plus penalty terms) with respect to $x$ and $z$, respectively. Next, we take the $x$-minimization step as an example to describe how it can be decomposed ($z$-minimization is similar).
As $x$-minimization focuses on the variable $x$ only, we can ignore terms related to $z$ in the augmented Lagrangian, such as the penalty terms from demand constraints. The resulting objective consists of three resource-related terms:
  • sum of per-resource allocation cost $f_{i}(x_{i*})$ from Equation 5;
  • sum of penalties for violating each resource constraint $R_{i}x_{i*}=r_{i}$ from Equation 2;
  • sum of penalties for violating $x=z$ from Equation 6.
The first two items are a sum of functions on per-resource allocation $x_{i*}$, while the third can also be rewritten as a sum of penalties for violating each $x_{i*}=z_{i*}$ (equivalent to $x=z$).
After rearranging, the optimization can be expressed as:
$\min_{x\in\mathcal{X}}\sum_{i}F_{i}(x_{i*}),$
where $F_{i}(x_{i*})$ consists of allocation cost, resource constraint penalty, and penalty for violating $x_{i*}=z_{i*}$. As each variable $x_{i*}$ is mutually exclusive and independent, the above problem is equal to optimizing per-resource allocations in parallel:
$\min_{x_{i*}\in\mathcal{X}_{i*}}F_{i}(x_{i*}),\quad\forall i,$
thereby decomposing the allocation of $x$ with $n\times m$ variables into $n$ per-resource allocations of $x_{i*}$, each with $m$ variables. To solve them, Zeal uses an off-the-shelf solver (cvxpy in our implementation) to find (exact or approximate) solutions while making liberal use of all available CPU cores.
Variables Objective
Boolean Integer Float Linear Convex
RDC[55]
SkyPilot[64]
ARROW[67], FlexWAN [39]
Shoofly[51]
PODP[7], RAS[46], Skyplane[26], Oort[34], TACCL[49], Shard Manager[37], Zeta[65], CASCARA[50], Sia[27], POP[43]
NetHint[14], Gavel[44], Teal[63], ONEWAN[32], BLASTSHIELD[31], NCFlow[1], Cerebro[66], DOTE[47], Soroush [42], POP[43]
PCF[29], Electricity Pricing[59], POP[43]
Table 1: Real-world resource allocation problems in recent literature.
4 Generality and Limitations
In this section, we start with the generality of Zeal, and then discuss its limitations, including scenarios with limited parallelism and chances of suboptimal solutions.
4.1 Generality
Zeal can be applied to a wide range of resource allocation problems. Table 1 presents a summary of real-world resource allocation problems in the recent literature, mostly from OSDI, SOSP, NSDI, and SIGCOMM in the past four years. Next, we explain why Zeal is applicable.
  • Variables.  Inheriting from ADMM, Zeal inherently supports real-number (float) allocation variables. Empirical evidence[53] and theoretical proofs [56, 38, 60] on ADMM also indicate that Zeal can effectively handle boolean and integer variables by projecting real-number solutions to desired data types over iterations.
  • Objective.  Zeal requires a max or sum of convex utility functions over per-resource or per-demand allocations. Most objectives used in systems exhibit convexity, such as the weighted allocation sum [44, 63], the logarithmic resource utility [3], and the quadratic allocation cost [64, 59].
  • Constraints.  Zeal assumes the linearity of resource and demand constraints, which holds in all our surveyed problems, e.g., the sum of demands allocated to a resource must not exceed its capacity. The constraints in these problems are per resource or per demand as well. Even under a constraint on multiple resources or demands, Zeal can be relaxed to groups of resources or demands. Zeal bundles these groups together into per-resource-group or per-demand-group allocations during decomposition.
Unlike POP [43], Zeal works with any workload, not only “granular” ones. In POP, the demands in each decomposed problem can access only a subset of resources, assuming each demand only requests a small fraction of interchangeable resources such that a resource subset suffices. This granular assumption can be unrealistic in practice, due to uneven demand distributions [28, 57] or limited resources that each demand is allowed to utilize [58, 33]. In contrast, each decomposed problem in Zeal can still access all the resources or demands; thus, the granular assumption is no longer required, affirming Zeal’s generality to various workload types.
4.2 Limitations
While Zeal applies to a variety of resource allocation problems, it might not always be the most suitable solution for every allocation problem.
Limited parallelism.  Zeal achieves parallelism by decomposing the full problem into per-resource and per-demand allocations. Therefore, when we cannot fully decompose the problem, Zeal will experience limited parallelism, although this does not happen in our survey (Table 1).
  • Non-separable objectives.  When the objective function is not a sum or max over allocation qualities per resource or per demand, we must bundle per-resource or per-demand allocations whose utilities stick together during decomposition. As a result, the partially decomposed problem will achieve limited parallelism.
  • Non-separable constraints.  When a resource constraint restricts the allocation across multiple resources, we must bundle the allocations for these resources during decomposition, which leads to limited parallelism. This also applies to demand constraints.
Suboptimal solutions.  It is possible for Zeal not to attain an ideal solution when the allocation problem deviates from the standard ADMM formulation.
  • Non-convex problems.  ADMM can solve non-convex optimization problems [53, 62, 68] empirically, providing theoretical guarantees only under specific assumptions[56, 38, 60]. As a result, Zeal can output suboptimal solutions in certain non-convex cases.
  • Nonlinear constraints.  When a constraint is nonlinear, we replace it by adding an indicator function to the objective. The indicator function is infinity when the constraint is violated or otherwise zero. However, this may alter the convexity of the objective function, resulting in the above non-convex problems.
  • Higher allocation dimension.  The allocation of resources to demands is encoded in a two-dimensional matrix $x$. Correspondingly, we separate the problem into two ADMM blocks for resources and demands. Introducing more dimensions, such as a temporal dimension [36], leads to more ADMM blocks and risks convergence failure [13].
5 Case Studies
In this section, we will present case studies on how Zeal can be applied to three resource allocation problems: traffic engineering, cluster scheduling, and load balancing, as also studied in POP [43].
5.1 Traffic Engineering
The traffic engineering problems [24, 54, 43, 42, 63] allocate traffic demand between datacenter nodes onto network links in a WAN topology.
Maximize total flow.  The problem focuses on maximizing the total flow allocation.
  • Variables.  The flow allocation from datacenter node pair $(s,t)$ onto network link $(u,v)$ is denoted as $x_{(u,v)(s,t)}\geq 0$. In common path-based traffic engineering, flows between node pair $(s,t)$ are allocated over links along pre-configured paths $P_{(s,t)}$ only.
  • Objective.  For a node pair $(s,t)$, the inflow it incurs to node $v$ is computed as $\text{inflow}(v,(s,t))=\sum_{(u,v)\in P_{s,t}}x_{(u,v)(s,t)}$ and outflow as $\text{outflow}(v,(s,t))=\sum_{(v,u)\in P_{s,t}}x_{(v,u)(s,t)}$. The total flow allocated for the node pair $(s,t)$ can be represented as $\text{inflow}(t,(s,t))$, which denotes the total inflow toward the destination $t$. Therefore, the objective of maximizing total flows between all node pairs can be expressed as the sum:
    $\operatorname*{Maximize\leavevmode\nobreak\ }_{x}\sum_{(s,t)}\text{inflow}(t,(s,t)).$
  • Resource constraints.  For each link $(u,v)$, the total flow allocation from all node pairs must not exceed its link capacity $c_{u,v}$:
    $\sum_{(s,t)}x_{(u,v)(s,t)}\leq c_{u,v},\quad\forall(u,v)\in E.$
  • Demand constraints.  For each node pair $(s,t)$, the total flow allocated cannot exceed the total demand. Additionally, for all nodes other than $s$ and $t$, the inflow and outflow must be equal, ensuring no flow “black hole” along the paths:
    $\begin{array}[]{l}0\leq\text{inflow}(t,(s,t))\leq d_{s,t},\\ \text{inflow}(v,(s,t))=\text{outflow}(v,(s,t)),\forall v\neq s,t,\\ \end{array}\quad\forall(s,t).$
Min-max link utilization.  The problem of min-max link utilization aims to allocate demands fairly over network links. The problem shares the same variables and linear demand constraints as the max total flow problem.
  • Objective.  The link utilization of a link $(u,v)$ is the ratio between the total flow allocated on that link and link capacity. Our objective is to minimize the maximum utilization of each link:
    $\operatorname*{Minimize\leavevmode\nobreak\ }_{x}\max_{(u,v)\in E}\frac{\sum_{(s,t)}x_{(u,v)(s,t)}}{c_{u,v}}.$
Using Zeal.  The traffic engineering problem aligns the separable structure for Zeal: the objectives are the sum of allocation quality per demand or the maximum of allocation quality per link, respectively. Linear constraints are applied to each demand and each link. Therefore, Zeal can decompose the problem into $|E|$ per-link allocations and $|V|^{2}$ per-demand allocations for parallelism.
5.2 Cluster Scheduling
Cluster scheduling problems [61, 44, 43, 42] involves allocating jobs to computing resources (e.g., CPUs, GPUs) in heterogeneous clusters.
Max-min allocation.  This problem aims to distribute the heterogeneous cluster resources fairly among jobs based on their effective throughput.
  • Variables.  With jobs time-sliced onto the available heterogeneous computing resources [44, 43], the allocation is a matrix $x\in[0,1]^{n\times m}$ where $x_{ij}$ is the fraction of time each job $j$ spent on each computing resource type $i$.
  • Objective.  Given an allocation matrix $x$, we adopt a normalized effective throughput of job $j$ from POP [41], which is a linear function on $x_{*j}$. The objective is to maximize the minimum allocation of normalized effective throughput:
    $\operatorname*{Maximize\leavevmode\nobreak\ }_{x}\min_{j}\leavevmode\nobreak\ \text{throughput}(j,x_{*j}).$
  • Resource constraints.  For each computing resource type $i$, the total time allocation cannot exceed the number of that computing resource type:
    $\sum_{j}x_{ij}z_{j}\leq\text{num\_workers}_{i},\quad\forall i,$
    where $z_{j}$ is the computing resources requested by job $j$.
  • Demand constraints.  For each job $j$, The sum of all fractions of time must be within 1:
    $\sum_{i}x_{ij}\leq 1,\quad\forall j.$
Proportional fairness.  The proportional fairness problem maximizes overall resource utilization while ensuring minimum service for each job. The problem shares the same variables, resource constraints, and capacity constraints.
  • Objective.  Proportional fairness aims to maximize the sum of log utilities of each job:
    $\operatorname*{Maximize\leavevmode\nobreak\ }_{x}\sum_{j}\log(\text{throughput}(j,x_{*j})).$
Using Zeal.  The cluster scheduling exhibits the separable structure required by Zeal: The objective function is a minimum of allocation quality per job and a sum of allocation quality per job, respectively. Additionally, linear constraints are applied to each job and each computing resource. Thus, Zeal can decompose the problem into $n$ per-job allocations and $m$ per-demand allocations.
5.3 Load Balancing
Load balancing problems [16, 48, 52, 43] allocate data shards among servers to scale out query loads in a distributed store.
Minimize shard movements.  This problem aims to minimize shard movements across servers during load changes while keeping the load nearly balanced on each server.
  • Variables.  The allocation is a matrix $x\in[0,1]^{n\times m}$, where $x_{ij}$ is the fraction of data shard $j$ allocated to server $i$. Additionally, a binary matrix $x^{\prime}$ captures shard placement, where $x^{\prime}_{ij}=1$ if the data shard $j$ is on the server $i$, i.e., $x_{ij}>0$; otherwise $x_{ij}=0$.
  • Objective.  The movement of data shard $j$ allocated to server $i$ can be represent as $(1-T_{ij})x^{\prime}_{ij}f_{i}$, where $T_{ij}$ denotes the initial shard placement and $f_{j}$ denotes the memory footprint of data shard $j$. The objective of minimizing total shard movement can be expressed as the sum:
    $\operatorname*{Minimize\leavevmode\nobreak\ }_{x}\sum_{j}\sum_{i}(1-T_{ij})x^{\prime}_{ij}f_{i}.$
  • Resource constraints.  For each server $i$, the total query load must be close to the average query load $L$; The total memory usage must be within memory capacity:
    $\begin{array}[]{l}L-\epsilon\leq\sum_{j}x_{ij}l_{j}\leq L+\epsilon,\\ \sum_{j}x^{\prime}_{ij}f_{j}\leq\text{memory}_{i},\end{array}\quad\forall i,$
    where $l_{j}$ is the query load on data shard $j$.
  • Demand constraints.  Each data shard $j$ must allocate all the query load among servers:
    $\sum_{i}x_{ij}=1,\quad\forall j.$
Using Zeal.  Given the separable structure, where the objective is the sum of allocation quality per data shard and linear constraints apply to each server, and each shard, Zeal can decompose the problem into $n$ per-server allocations and $m$ per-shard allocations.
6 Implementation of Zeal
We implement Zeal as a Python package. After publication, we will upload Zeal to PyPI for easy installation (pip install zeal).
Listing 1 showcases a simple resource allocation problem solved using our Zeal package. In line 4, we initially create a $N\times M$ non-negative allocation variable x, where x[i,j] represents the allocation of resource i to demand j. We then create a constraint for each resource in line 7–8 and a constraint for each demand in line 9–10, as well as an objective of minimizing the sum of x in line 13. Finally, we construct the problem in line 16–17 and solve it in line 18.
1import zeal
2
3# Create allocation variables.
4x = zeal.Variable((N, M), nonneg=True)
5
6# Create the constraints.
7resource_constrs = [
8 x[i,:].sum() <= 1 for i in range(N)]
9demand_constrs = [
10 x[:,j].sum() <= 1 for j in range(M)]
11
12# Create an objective.
13obj = zeal.Minimize(x.sum())
14
15# Construct and solve the problem.
16prob = zeal.Problem(
17 obj, resource_constrs, demand_constrs)
18prob.solve(num_cpus=64)
Listing 1: Resource allocation example with Zeal package.
Zeal borrows the interface from cvxpy [17, 4]. Zeal inherits most of cvxpy’s methods, e.g. Variable(.) in line 4, Minimize(.) in line 13. One difference from cvxpy is that users must specify resource constraints and demands constraints separately (line 16–17) when initializing a problem. As a parallel solution, Zeal allows users to specify the number of CPU cores (line 18); if not specified, Zeal uses all available CPU cores.
Behind Zeal’s interface, the resource allocation problem is solved in the following three steps.
  • Problem parsing.  Inequality constraints are converted into equality constraints with a non-negative slack variable s., e.g., inequality x[0,:].sum() <= 1 can be converted to equality x[0,:].sum() + s == 1. The slack variable can be regarded as an additional part of the problem variables and be optimized together.
  • Problem building.  Resource constraints are organized into independent per-resource subgroups without shared variables, while demand constraints are similarly organized into separate independent per-demand groups. For each subgroup, a per-resource or per-demand problem is built using cvxpy, with the new objective formulated based on the augmented Lagrangian method outlined in § 3.2.
  • Problem solving.  Per-resource or per-demand problems in cvxpy are solved in parallel across multiple cores. Between ADMM iterations, only the parameters are updated, avoiding the need to rebuild the entire cvxpy problem. Similarly, for the same problem with varying resources and demands, only the corresponding parameters are updated.
We use Ray [40] to parallelize our computation across multiple cores. Python has limited parallelism in its multithreading implementation due to the notorious global interpreter lock (GIL). Instead, we choose Ray, which provides a wrapper for multiple Python interpreter processes, so as to avoid GIL. We also benefit from Ray’s production-grade implementation of inter-process coordination mechanisms to avoid building them from scratch using Python’s built-in low-level multiprocessing library.
7 Evaluation
We answer the following questions in the evaluation:
  1. 1.
    How does Zeal compare with state-of-the-art approaches in allocation quality and computation time? (§7.1)
  2. 2.
    How do Zeal and state-of-the-art approaches react to changes in problem granularity, temporal fluctuation, spatial redistribution, and link failures? (§7.2)
  3. 3.
    What is the contribution of each design in Zeal? (§7.3)
We focus on four approaches throughout our evaluation:
  • Exact Sol.  Exact Sol solves the original resource allocation problem using commercial optimization solvers. We reuse the implementation from POP, which employs Gurobi [22], cvxpy [17, 4], and CPLEX [15] for traffic engineering, cluster engineering, and load balancing, respectively.
  • POP.  POP-$k$ [43] randomly splits the problem into $k$ smaller subproblems, applies commercial solvers in each, and coalesces the resulting $k$ sub-allocations into a global allocation. However, POP only simulates the parallel execution by solving subproblems sequentially, collecting the solving time of each, and calculating the parallel solving time mathematically.
  • Zeal.  Zeal is our truly parallel implementation, with details described in Section 6.
  • Zeal.  To facilitate a fair comparison with POP, we create a variant of Zeal, denoted as Zeal, following POP-style simulation where subproblems are solved sequentially and a parallel solving time is calculated mathematically.
We also consider domain-specific approaches apart from these four approaches. In traffic engineering, we evaluate a demand-pinning approach (“Pinning” for short) similar to a prior work [41], with the top 10% of demands allocated using optimization engines and the remaining assigned to the shortest paths. We also evaluate Teal [63], a learning-accelerated traffic engineering approach with massive parallelism achieved on the GPU. In cluster scheduling and load balancing, we evaluate the greedy heuristic algorithms, Gandiva [61] and E-Store [52], respectively.
We compare these approaches based on their allocation quality and computation times. The computation time is measured on 64 CPU cores (2$\times$ Intel Xeon Gold 6142); an additional GPU (Nvidia Titan RTX) is made available for Teal.
7.1 Zeal vs. the State of the Art
In this section, we compare Zeal against state-of-the-art approaches on resource allocation problems from Section 5.
7.1.1 Traffic Engineering
Figure 3: Comparison for the problem of maximizing total flow in traffic engineering.
We adopt the largest network in Teal’s dataset for traffic engineering problems. This 1,739-node network is adapted from an AS-level internet topology [12] for WAN purposes, with the traffic data collected from a global cloud provider and then mapped into the large network.
Maximize total flow.  Figure 3 compares Zeal to state-of-the-art methods in maximizing total flow. Zeal achieves 90.8% satisfied demand within 30 seconds and 92% within 60 seconds, demonstrating a superior trade-off between satisfied demand and computation time compared to Exact Sol, POP, and Pinning. While POP-4 also reaches 92% satisfied demand, its average computation time is significantly longer at 1658 seconds. POP-16 and POP-64 improve parallelism by dividing the workload into more subproblems, reducing computation times to 456 seconds and 380 seconds, respectively. Note that the speedup of POP-64 is limited compared to POP-16 as the number of cores per subproblem drops from 4 to 1. This increased parallelism also comes at a cost, with satisfied demand decreasing to 87.4% for POP-16 and 81.6% for POP-64. Pinning focuses on top demands to reduce problem size but remains constrained by the sequential nature of its optimization solvers. As a result, it achieves a modest speedup, requiring 149 seconds to reach 87.3% satisfied demand. Teal, a learning-accelerated approach, leverages massive GPU parallelism to achieve 89% satisfied demand in just 1 second. However, it relies on a highly customized machine-learning framework, demanding significant human effort for design and training.
Zeal, representing the same algorithm under ideal conditions as POP, achieves a satisfied demand that is 8% higher than Zeal within 10 seconds in Figure 3 and consistently delivers superior allocation quality across the other figures. The reason is that Zeal executes each iteration more quickly, allowing it to complete more iterations within a given time, resulting in higher allocation quality. The difference in iteration time between Zeal and Zeal arises from three factors, as discussed in more detail in § 7.3: (1) Zeal measures the end-to-end time, including problem-solving, compilation, solution unpacking, and other associated tasks, while Zeal only takes the problem-solving time into account. (2) Zeal’s parallel implementation (§ 6) incurs overhead from cache contention, which slows down the problem-solving process. (3) In Zeal, each subproblem is statically pre-assigned to one of the processes, making it susceptible to stragglers. In contrast, Zeal assumes perfect dynamic scheduling.
Figure 4: Comparison for the problem of minimizing max link utilization in traffic engineering.
Min-max link utilization.  Figure 4 compares Zeal with other approaches in minimizing maximum link utilization. Zeal achieves a low maximum link utilization of 1.67 within 10 seconds and 1.71 within 5 seconds. Exact Sol reaches 1.63 with an average computation time of 35 seconds, and it is faster than itself in maximizing total flow because minimizing max link utilization is "easier" and requires fewer iterations. Pinning takes 42.5 seconds, making it slower than Zeal and slightly slower than Exact Sol, as it requires additional time to rebuild the problem when the top 10% of node pairs change. POP’s variants, POP-4, POP-16, and POP-64, achieve a maximum link utilization of 1.70, 1.77, and 1.95 with 9, 2, and 0.4 seconds on average, respectively. Although POP-16 is faster than Zeal, a fairer comparison between POP-16 and Zeal, both simulating parallelism, reveals that Zeal achieves a comparable computation time to POP-16. The domain-specific solution Teal takes only 0.3 seconds with massive parallelism from GPU with a maximum link utilization of 1.74.
7.1.2 Cluster Scheduling
We evaluate a large cluster scheduling problem of allocating jobs to 5000 GPU/CPUs of 196 different types. We collect real throughput from two computing benchmarks [25, 35]. The number of GPU/CPUs for each type and the required number in each job is generated based on the distribution observed in traces from real GPU-sharing clusters [58]. To reduce simulation time, we capture snapshots through the scheduling process rather than executing the entire simulation from start to finish.
Figure 5: Comparison for the min-max allocation problem in cluster scheduling.
Max-min allocation.  Figure 5 compares Zeal against other approaches with the objective of max-min allocation. Zeal achieves a near-optimal max-min allocation of 0.97 within 5 seconds and 0.98 within 10 seconds, while Zeal reaches 0.98 within 2 seconds. Due to their sequential nature, Exact Sol and Gandiva encounter a slow computation time of 86 and 22 seconds. The greedy heuristic Gandiva even experiences a 0.56 low max-min allocation. POP-4 deliverers a comparable time-accuracy tradeoff as Zeal but is 3.4$\times$ slower than Zeal, which is simulated in the same way as POP did for a fair comparison. A faster POP variant, POP-16, exhibits a lower max-min allocation of 0.94, as it becomes more difficult for jobs to assess the ideal GPU/CPU among the limited selection in smaller subproblems.
Figure 6: Comparison for the problem of maximizing proportional fairness in cluster scheduling.
(a) Granularity changes.
(b) Temporal fluctuation.
(c) Spatial redistribution.
Figure 7: Normalized satisfied demand in traffic engineering against granularity, temporal, and spatial changes.
Figure 8: Comparison for the problem of minimizing shard movements in load balancing.
Proportional fairness.  Figure 6 compares Zeal against other approaches in maximizing proportional fairness. As the optimization problem becomes more difficult with the convex logarithm function in the objective, both Zeal and Zeal achieve a greater speedup than POP by thoroughly breaking down the problem into smaller and easier pieces. Zeal and Zeal reach a proportional fairness score of over 1,000 in 300 seconds and 100 seconds, respectively. In contrast, POP-4 takes 1,825 seconds to reach the same score, while POP-16 and POP-64 remain below 900 despite their shorter computation times. Exact Sol, though accurate, is prohibitively slow, requiring 12,432 seconds.
7.1.3 Load Balancing
We evaluate a load balancing problem of assigning 2048 data shards to 256 servers. In each round, we compute a new allocation of shards to servers given the new query loads [43].
Minimize shard movements.  Figure 8 compares Zeal with state-of-the-art approaches in minimizing shard movements. Zeal achieves an average of 24.1 shard movements in 15 seconds, slightly better than POP-4’s 25.4 shard movements within 16.9 seconds. In faster scenarios, Zeal achieves 35.1 shard movements in 2 seconds, while POP-4 achieves slightly fewer movements of 32.2 in 1.7 seconds. In a more fair comparison with POP, Zeal consistently costs 15–18% fewer shard movements than POP given a similar time limit. Exact Sol achieves the minimum shard movements of 13.2 but is significantly slower, taking an average of 1,572 seconds due to the NP-hardness of solving the mixed-integer linear programming problem inherent to load balancing. On the other hand, the greedy heuristic represents the opposite end of the time-accuracy trade-off, completing in just 25 milliseconds but requiring 50 shard movements.
(a) Speedup when varying CPU cores.
(b) Convergence rate.
(c) Comparison to more optimization methods.
Figure 9: Micro-benchmarks with the objective of maximizing total flow in traffic engineering.
Figure 10: Satisfied demand of traffic engineering under zero, 50, 100, or 200 link failures after recomputing flow allocation.
7.2 Robustness
We evaluate the robustness of Zeal and other baselines against changes in problem granularity, temporal fluctuations, spatial redistribution, and link failures. In all these evaluations, Zeal is run for 30 seconds. To mitigate the impact of the changing resources and demands, we normalize the satisfied demand by the optimal satisfied demand from Exact Sol.
Changes of problem granularity.  The granular problem required by POP, where resources are mutually interchangeable, is impractical in reality. In traffic engineering, for example, demands are constrained to capacity resources on pre-configured paths. Similarly, in cluster scheduling, 33% of GPU tasks in production clusters are limited to specific GPU types [58]. Using traffic engineering as an example, we evaluate all approaches across varying levels of problem granularity, as depicted in Figure 7(a). To quantify resource interchangeability, we use the mean edge betweenness centrality [10, 11], which measures the average percentage of demands served by a given edge.
As the mean edge betweenness centrality decreases, all approaches experience a decline in normalized satisfied demand, ranging from 0.8–2.1% for Zeal, 2.8–4.2% for Teal, 2.5–5.6% for Pinning, and 1–12.3% for POP. However, the maximum decrease in POP is 5.9$\times$ greater than Zeal because POP’s subproblems struggle to find alternative resources within their restricted subsets, whereas Zeal ’s subproblems retain access to the full set of resources. Teal exhibits a slight 1.3% increase in satisfied demand when the mean edge betweenness centrality is minimal. This is because the reduced edge selection simplifies allocation decisions.
Temporal fluctuation.  We introduce temporal fluctuations [63] to the traffic matrices for robustness assessment. For each demand, we calculate the variance $\sigma^{2}$ in its changes between consecutive time slots and create a new normal distribution $N(0,k\sigma^{2})$. We choose the values of $k$ to be 2, 5, 10, and 20. Next, we randomly draw a sample from the newly created normal distribution and add it to each demand in every time slot, simulating temporal fluctuations.
Figure 7(b) demonstrates that Zeal effectively manages temporal fluctuations, with changes in normalized satisfied demand limited to 1.3–3.5%. In contrast, Teal suffers a significant decrease of 12.8%, as it lacks prior exposure to the distribution of these fluctuating demands during training. Pinning and POP experience variations in normalized satisfied demand ranging from 0.6–3% and 1–1.6%, respectively. This is because randomness may occasionally favor splitting a problem into subproblems or focusing on top demands exclusively, while at other times, it does not.
Spatial redistribution.  To evaluate robustness against spatial redistribution, we reassign traffic demands among node pairs. Specifically, we adjust the top 10% of demands, which initially account for 88.4% of the total volume, so that they take up 80%, 60%, 40%, and 20% of the total volume instead. Figure 7(c) shows that Zeal consistently fulfills the highest demand across all spatial distributions. Pinning’s performance decreases by 8.3%, as its strategy of focusing only on top demands relies on the presence of a heavy-tailed demand distribution. Teal has a slight decrease up to 2.4% due to unseen demand distribution in training. This is a relatively smaller decrease than the temporal fluctuation, as Teal can still handle simple, evenly distributed traffic demands. POP has a slight increase of 1.4% as demands are more granular, where no demand requires a significant fraction of the total resources.
Link failures.  In practice, large-scale inter-datacenter link failures are exceedingly rare. Even in the event of fiber link failures, they typically do not result in capacity loss on an inter-datacenter link, unless caused by a fiber cut [51]. As a stress test, however, we assess extreme failure scenarios described in previous work [67] and deliberately introduce 50, 100, and 200 link failures. Figure 10 illustrates the offline satisfied demand following these link failures, and we observe a consistent trend across different link failure scenarios across all approaches. Specifically, all approaches drop 0.6–1.1% under 50 link failures, 1.5–2.1% under 100 link failures, and 3–4.2% under 200 link failures. The reason is that these link failures constitute only a minor portion of the total 8,558 links, allowing all approaches to exhibit robustness given adequate time for re-computations.
7.3 Micro-benchmarks
We conduct micro-benchmarks to evaluate the impact of Zeal’s design components, with all evaluations performed in maximizing total flow in traffic engineering.
Speedup when varying CPU cores.  Figure 9(a) evaluates the speedup of Zeal, Zeal, and Exact Sol with 1, 4, 16, and 64 CPU cores. Zeal achieves 61.7$\times$ speedup under 64 cores, where the slight gap occurs as we cannot exactly divide the total computation time evenly across 64 cores. Zeal also achieves an almost linear speedup when the number of CPU cores grows from 1 to 16. However, this linearity diminishes with 64 cores, where Zeal only achieves 18.2$\times$ speedup. This is because utilizing all 64 cores for parallelism incurs overhead from cache contention and stragglers, as discussed in Section 7.1. In contrast to the linear speedup of Zeal and Zeal, the speedup of Exact Sol is sublinear and marginal, achieving only 3.4$\times$ speeds with 64 CPU cores. This is because of commercial solvers’ sequential nature, e.g., the simplex method [45] in linear programming requires thousands to millions of small steps to reach the optimal solution.
Convergence rate.  Both Teal and Zeal employ ADMM for parallel optimization, but their approaches differ. Teal’s ADMM adapts its computations based on matrix multiplication to leverage GPU calculation capabilities. Figure 9(b) illustrates the convergence rates of Teal and Zeal under various initialization. Zeal takes 30 seconds to surpass 90% satisfied demand. However, when initialized with Teal’s coarse neural network solution, Zeal achieves this within 20 seconds. Conversely, with naive initialization, Zeal requires more than 60 seconds. Teal, benefiting from the massive parallelism of GPUs, reaches 90% satisfied demand within 10 seconds. However, its adapted ADMM limits its performance, capping its satisfied demand at 91% at 60 seconds compared to Zeal’s 92%.
Comparison to more optimization methods.  Figure 9(c) investigates more constrained optimization methods other than Zeal’s ADMM method in maximizing the total flow in traffic engineering. Section 7.1 has covered the computation time of the optimal Exact sol, the suboptimal POP-64, and the near-optimal Zeal. As Zeal replaces constraints with additional terms in the objective function, we compare similar methods, including the penalty method and the augmented Lagrangian method. The augmented Lagrangian method is the same as in Zeal’s ADMM but without decoupling and decomposition, and thus it is over 3$\times$ slower than Zeal in reaching a satisfied demand of over 90%. The penalty method, which gradually increases penalty parameters from 0 to infinity to enforce the converted constraints in the objective function, converges slowly and is over 30$\times$ slower than Zeal in reaching a satisfied demand of over 90%.
8 Related Work
Resource allocation problems.  Multi-tenant systems often address resource allocation challenges. Section 5 examines three prominent examples: cluster scheduling [46, 27, 44, 61, 64] allocates jobs to CPUs or GPUs in heterogeneous clusters, traffic engineering [24, 1, 67, 50, 31] allocates traffic demands to network link capacities in wide-area networks, and load balancing [16, 48, 52] allocates query loads to servers in distributed databases. Beyond these problems, Skyplane [26] allocates cloud resources (e.g., bandwidth and TCP connections) to overlay paths that traverse intermediate cloud regions, minimizing the cost of inter-region bulk transfers. Zeta [65] allocates traffic forwarding capability from gateway clusters to virtual private clouds for scalable and robust east-west communication in large-scale clouds. Shoofly [51] allocates light wavelengths of optical fiber to network shortcuts abstracted from optical bypasses, in order to minimize the hardware costs of provisioning long-haul capacity.
Optimization algorithms.  Various constrained optimization algorithms have been employed to solve resource allocation problems, depending on the variable types, objectives, and constraints. For linear programming problems like traffic engineering, the Gurobi solver [22] utilizes two primary methods. The simplex method [45] iteratively progresses along the boundaries of the feasible region toward the optimal solution, while the barrier method [30] iteratively approaches the optimal solution from within the feasible region. In mixed-integer linear programming like load balancing, the branch-and-bound method [8] divides the problem into smaller subproblems, or branches, and then eliminates certain branches based on bounds on the optimal solution at each iteration. For more complex optimization problems, such as cone programming when maximizing proportional fairness in cluster scheduling, the ECOS algorithm [18] adopts a similar iterative process as the barrier methods, with search directions found by a symmetric indefinite KKT system. All these constrained optimization algorithms operate sequentially and iteratively, and thus introduce bottlenecks in solving time.
Decomposition for scalability.  As resource allocation problems grow, efforts have been made to parallelize the iterative process in optimization algorithms for better scalability. NCFlow [1] accelerates traffic engineering by partitioning the network into clusters and concurrently solving the flow allocation problem in each cluster. POP [43] targets “granular” allocation problems where resources are interchangeable. It randomly splits resources and demands into subsets and allocates each subset in parallel. Soroush [42] generalizes the classical waterfilling algorithm for max-min fair allocation and presents parallelizable combinatorial algorithms with fairness guarantees. Teal [63] models traffic engineering with neural networks to benefit from the massive parallelism in GPU. Eason et al. [19] employ the Benders decomposition algorithm to parallelize the design of a specialized cross-layer backbone network, offering performance guarantees for a particular allocation problem.
9 Conclusion
In this paper, we scale real-world resource allocation problems from a fresh perspective, observing that most of them exhibit an inherent separable structure. This key insight allows us to systematically decouple resource and demand constraints and decompose the original optimization problem into independent and parallel subproblems for allocating individual resources and demands. We implement this approach, Zeal, as a library to augment a widely-used, open-source solver (cvxpy), offering a push-button solution that accelerates large-scale resource allocation. Through extensive evaluation across diverse resource allocation tasks, we demonstrate that Zeal significantly improves scalability while closely approximating the quality of exact solutions.
Appendix A Mathematical Representation of Zeal
A resource allocation problem can be written as
$\min_{x\in\mathcal{X}}\quad$ $\sum_{i}f_{i}(x_{i*})+\sum_{j}g_{j}(x_{*j}),$ (1)
$s.t.\quad$ $R_{i}x_{i*}=r_{i},\quad\forall i,$ (2)
$D_{j}x_{*j}=d_{j},\quad\forall j,$ (3)
where Equation 1 is the objective, and Equation 2 and Equation 3 are resource constraints and demand constraints, respectively.
Our first step decouples resource and demand constraints. We create a copy of the allocation variable $x$, denoted as $z$, and replace the variable $x$ in demand constraints (Equation 3) as well as the demand-related portion of the objective function (Equation 1) with the variable $z$. We also add a new constraint $x=z$ (Equation 6) to keep $x$ and $z$ equal.
$\min_{x\in\mathcal{X}}\quad$ $\sum_{i}f_{i}(x_{i*})+\sum_{j}g_{j}(z_{*j}),$ (1)
$s.t.\quad$ $R_{i}x_{i*}=r_{i},\quad\forall i,$ (2)
$D_{j}z_{*j}=d_{j},\quad\forall j,$ (3)
$x-z=0.$ (6)
We rewrite our problem in the Augmented Lagrangian to move the constraints to the objective function:
$\begin{split}&\mathcal{L}_{\rho}(x,z,\alpha,\beta,\lambda)=\sum_{i}f_{i}(x_{i*})+\sum_{j}g_{j}(z_{*j})\\ &+\frac{\rho}{2}\Bigl{(}\sum_{i}||R_{i}x_{i*}-r_{i}+\alpha_{i}||_{2}^{2}+\sum_{j}||D_{j}z_{*j}-d_{j}+\beta_{j}||_{2}^{2}\\ &+||x-z+\lambda||_{F}^{2}\Bigr{)}-\frac{\rho}{2}(||\alpha||_{2}^{2}+||\beta||_{2}^{2}+||\lambda||_{F}^{2}).\end{split}$
ADMM allows partial optimization alternatively on $x$ and $z$:
$x^{(k+1)}$ $:=\operatorname*{arg\,min}_{x\in\mathcal{X}}\mathcal{L}_{\rho}(x,z^{(k)},\alpha^{(k)},\beta^{(k)},\lambda^{(k)})$ (7)
$z^{(k+1)}$ $:=\operatorname*{arg\,min}_{z}\mathcal{L}_{\rho}(x^{(k+1)},z,\alpha^{(k)},\beta^{(k)},\lambda^{(k)})$ (8)
$\alpha_{i}^{(k+1)}$ $:=\alpha_{i}^{(k)}+R_{i}x^{(k+1)}_{i*}-r_{i},\forall i$
$\beta_{j}^{(k+1)}$ $:=\beta_{j}^{(k)}+D_{j}z^{(k+1)}_{*j}-d_{j},\forall j$
$\lambda^{(k+1)}$ $:=\lambda^{(k)}+x^{(k+1)}-z^{(k+1)}$
Our second step further decomposes the partial optimization into optimizing each row in $x$ and each column in $z$. We rewrite Equation 7 as
$x^{(k+1)}=$ $\operatorname*{arg\,min}_{x\in\mathcal{X}}\mathcal{L}_{\rho}(x,z^{(k)},\alpha^{(k)},\beta^{(k)},\lambda^{(k)})$ (7)
$=$ $\operatorname*{arg\,min}_{x\in\mathcal{X}}\sum_{i}f_{i}(x_{i*})+\frac{\rho}{2}\sum_{i}\big{\|}R_{i}x_{i*}-r_{i}+\alpha_{i}^{(k)}\big{\|}_{2}^{2}$
$+\frac{\rho}{2}\big{\|}x-z^{(k)}+\lambda^{(k)}\big{\|}_{F}^{2}.$
We rearrange the parentheses in Equation 2 to make the variables in each parenthesis mutually exclusive:
$x^{(k+1)}=$ $\operatorname*{arg\,min}_{x\in\mathcal{X}}\sum_{i}\Bigl{(}f_{i}(x_{i*})+\frac{\rho}{2}\big{\|}R_{i}x_{i*}-r_{i}+\alpha_{i}^{(k)}\big{\|}_{2}^{2}$
$+\frac{\rho}{2}\big{\|}x_{i*}-z_{i*}^{(k)}+\lambda_{i*}^{(k)}\big{\|}_{2}^{2}\Bigr{)}.$
With the variable in each parenthesis mutually exclusive, we decompose to optimize per-resource allocation $x_{i*}$ individually:
$\begin{split}x_{i*}^{(k+1)}=&\operatorname*{arg\,min}_{x_{i*}\in\mathcal{X}_{i*}}\Bigl{(}f_{i}(x_{i*})+\frac{\rho}{2}\big{\|}R_{i}x_{i*}-r_{i}+\alpha_{i}^{(k)}\big{\|}_{2}^{2}\\ &+\frac{\rho}{2}\big{\|}x_{i*}-z_{i*}^{(k)}+\lambda_{i*}^{(k)}\big{\|}_{2}^{2}\Bigr{)}.\end{split}$
We follow a similar decomposing procedure for demand-constrained optimization on $z$:
$z^{(k+1)}=$ $\operatorname*{arg\,min}_{z}\mathcal{L}_{\rho}(x^{(k+1)},z,\alpha^{(k)},\beta^{(k)},\lambda^{(k)})$ (8)
$=$ $\operatorname*{arg\,min}_{z}\sum_{j}g_{j}(z_{*j})+\frac{\rho}{2}\sum_{j}\big{\|}D_{j}z_{*j}-d_{j}+\beta_{j}^{(k)}\big{\|}_{2}^{2}$
$+\frac{\rho}{2}\big{\|}x^{(k+1)}-z+\lambda^{(k)}\big{\|}_{F}^{2}$
$=$ $\operatorname*{arg\,min}_{z}\sum_{j}\Bigl{(}g_{j}(z_{*j})+\frac{\rho}{2}\big{\|}D_{j}z_{*j}-d_{j}+\beta_{j}^{(k)}\big{\|}_{2}^{2}$
$+\frac{\rho}{2}\big{\|}x^{(k+1)}_{*j}-z_{*j}+\lambda_{*j}^{(k)}\big{\|}_{2}^{2}\Bigr{)}.$
The optimization of $z$ can now be decomposed into optimization on per-demand allocation $z_{*j}$:
$\begin{split}z_{*j}^{(k+1)}=&\operatorname*{arg\,min}_{z_{*j}}\Bigl{(}g_{j}(z_{*j})+\frac{\rho}{2}\big{\|}D_{j}z_{*j}-d_{j}+\beta_{j}^{(k)}\big{\|}_{2}^{2}\\ &+\frac{\rho}{2}\big{\|}x^{(k+1)}_{*j}-z_{*j}+\lambda_{*j}^{(k)}\big{\|}_{2}^{2}\Bigr{)}.\end{split}$

### v2

@@
-Zeal: Rethinking Large-Scale Resource Allocation with “Decouple and Decompose”
+Decouple and Decompose: Scaling Resource Allocation with DeDe
 Abstract
@@
 In this work, we revisit real-world resource allocation problems and uncover a common underlying structure: a vast majority of these problems are inherently separable, i.e., they optimize the aggregate utility of individual resource and demand allocations, under separate constraints for each resource and each demand.
-Building on this insight, we develop Zeal, a general, scalable, and theoretically grounded framework for accelerating resource allocation through a “decouple and decompose” approach.
-Zeal systematically decouples entangled resource and demand constraints, thereby decomposing the overall optimization into alternating per-resource and per-demand allocations, which can then be solved efficiently and in parallel.
-We have implemented Zeal as a library extension to an open-source solver, maintaining a familiar user interface.
-Experimental results across three prominent resource allocation tasks—traffic engineering, cluster scheduling, and load balancing—demonstrate Zeal’s substantial speedups and robust allocation quality.
+Building on this insight, we develop DeDe, a general, scalable, and theoretically grounded framework for accelerating resource allocation through a decouple-and-decompose approach.
+DeDe systematically decouples entangled resource and demand constraints, thereby decomposing the overall optimization into alternating per-resource and per-demand allocations, which can then be solved efficiently and in parallel.
+We have implemented DeDe as a library extension to an open-source solver, maintaining a familiar user interface.
+Experimental results across three prominent resource allocation tasks—traffic engineering, cluster scheduling, and load balancing—demonstrate DeDe’s substantial speedups and robust allocation quality.
 1 Introduction
@@
 In this study, we re-examine real-world resource allocation problems through a different lens and aim to develop a general and scalable optimization framework without catering to specific problems or making workload-dependent assumptions.
-Our approach, Zeal, is theoretically grounded and enables fine-grained decomposition of the optimization problem, alternating between per-resource and per-demand allocations that can be solved in parallel (Figure 1).
-Figure 1: Overview of Zeal.
+Our approach, DeDe, is theoretically grounded and enables fine-grained decomposition of the optimization problem, alternating between per-resource and per-demand allocations that can be solved in parallel (Figure 1).
+Figure 1: Overview of DeDe.
 This decomposition is made possible by an intrinsic, workload-agnostic problem structure that we uncovered after surveying dozens of real-world resource allocation problems in recent literature (Table 1).
@@
 Ignoring the dependencies across subproblems and greedily solving each subproblem independently would fail to find feasible and high-quality allocations.
-Given the intertwined nature of resource and demand constraints, Zeal first decouples them as follows.
+Given the intertwined nature of resource and demand constraints, DeDe first decouples them as follows.
 Let $x$ denote the resource allocation matrix between $n$ resources and $m$ demands, where $x_{ij}$ represents the amount of resource $i$ assigned to demand $j$.
-The key mathematical trick in Zeal is to introduce a duplicate of the matrix $x$, denoted as $z$, along with an additional constraint $x=z$.
-Next, Zeal iteratively and alternately solves for $x$ subject only to resource constraints with $z$ fixed, and solves for $z$ subject only to demand constraints with $x$ fixed.
-In each iteration, Zeal further ensures that $x$ and $z$ do not deviate significantly from each other in order to satisfy $x=z$.
+The key mathematical trick in DeDe is to introduce a duplicate of the matrix $x$, denoted as $z$, along with an additional constraint $x=z$.
+Next, DeDe iteratively and alternately solves for $x$ subject only to resource constraints with $z$ fixed, and solves for $z$ subject only to demand constraints with $x$ fixed.
+In each iteration, DeDe further ensures that $x$ and $z$ do not deviate significantly from each other in order to satisfy $x=z$.
 This decoupling transformation lends itself to a classical constrained optimization framework called ADMM (Alternating Direction Method of Multipliers) [9].
-ADMM allows Zeal to embed each of the resource and demand constraints into the original objective as a penalty for constraint violation, thereby constructing augmented Lagrangians [23] as new optimization objectives.
+ADMM allows DeDe to embed each of the resource and demand constraints into the original objective as a penalty for constraint violation, thereby constructing augmented Lagrangians [23] as new optimization objectives.
 ADMM provides the theoretical foundation that enables alternating optimizations of these augmented Lagrangians with respect to $x$ and $z$, respectively.
-Next, Zeal further decomposes the above optimizations into per-resource and per-demand allocations.
+Next, DeDe further decomposes the above optimizations into per-resource and per-demand allocations.
 This is feasible owing to the separable structure we identified across resource allocation problems: the original objective is separable on each resource and demand variable, and so are the penalty terms converted by ADMM from constraints in the augmented Lagrangians.
@@
 These subproblems can be solved independently with massive parallelism using off-the-shelf solvers.
-Figure 2 illustrates the “decouple and decompose” workflow of Zeal.
-To demonstrate Zeal’s practicality, we have implemented Zeal as a Python package [^3] Zeal can be installed with pip install zeal after publication., built on the widely used open-source solver cvxpy [17, 4].
-The interface of Zeal is also designed to resemble cvxpy for ease of use (§6).
-Moreover, instead of simulating the parallel execution of subproblems on multiple CPU cores as previous studies did [1, 43], we have carefully optimized Zeal’s implementation using Ray [40], enabling efficient utilization of multiple cores without incurring significant overheads, such as cache contention.
-We evaluate Zeal in three prominent resource allocation problems—traffic engineering, cluster scheduling, and load balancing—and demonstrate its faster speed and higher allocation quality than the state of the art (§7).
-Compared with the best variant of POP [43] for each problem, Zeal achieves a 5.3% higher allocation quality and a 7.6$\times$ speedup in traffic engineering, 14% and 1.34$\times$ in cluster scheduling, and 5.3% and 1.13$\times$ in load balancing.
-Figure 2: Workflow of Zeal.
+Figure 2 illustrates the “decouple and decompose” workflow of DeDe.
+To demonstrate DeDe’s practicality, we have implemented it as a Python package, installable via pip install dede.
+DeDe is built on the widely used open-source solver cvxpy, and its interface is also designed to resemble cvxpy for ease of use (§6).
+Moreover, instead of simulating the parallel execution of subproblems on multiple CPU cores as previous studies did [1, 43], we have carefully optimized DeDe’s implementation using Ray [40], enabling efficient utilization of multiple cores without incurring significant overheads, such as cache contention.
+We evaluate DeDe in three prominent resource allocation problems—traffic engineering, cluster scheduling, and load balancing—and demonstrate its faster speed and higher allocation quality than the state of the art (§7).
+Compared with the best variant of POP [43] for each problem, DeDe achieves a 5.3% higher allocation quality and a 7.6$\times$ speedup in traffic engineering, 14% and 1.34$\times$ in cluster scheduling, and 5.3% and 1.13$\times$ in load balancing.
+Figure 2: Workflow of DeDe.
 2 Real-World Resource Allocation Problems
@@
 Next, we describe our solution to this challenge.
-3 Zeal: Decouple and Decompose
-We present Zeal, a general, scalable, and theoretically rooted framework for optimizing real-world resource allocation problems.
-Leveraging the separable structure (§2) that underpins a vast majority of these problems, Zeal enables massive parallelism with near-optimal solutions through a technique that we call “decouple and decompose.” A formal, mathematical representation of Zeal is included in Appendix A.
+3 DeDe: Decouple and Decompose
+We present DeDe, a general, scalable, and theoretically rooted framework for optimizing real-world resource allocation problems.
+Leveraging the separable structure (§2) that underpins a vast majority of these problems, DeDe enables massive parallelism with near-optimal solutions through a decouple-and-decompose technique.
+A formal, mathematical representation of DeDe is included in Appendix A.
 3.1 Decouple
@@
 thereby decomposing the allocation of $x$ with $n\times m$ variables into $n$ per-resource allocations of $x_{i*}$, each with $m$ variables.
-To solve them, Zeal uses an off-the-shelf solver (cvxpy in our implementation) to find (exact or approximate) solutions while making liberal use of all available CPU cores.
+To solve them, DeDe uses an off-the-shelf solver (cvxpy in our implementation) to find (exact or approximate) solutions while making liberal use of all available CPU cores.
 Variables Objective
@@
 4 Generality and Limitations
-In this section, we start with the generality of Zeal, and then discuss its limitations, including scenarios with limited parallelism and chances of suboptimal solutions.
+In this section, we start with the generality of DeDe, and then discuss its limitations, including scenarios with limited parallelism and chances of suboptimal solutions.
 4.1 Generality
-Zeal can be applied to a wide range of resource allocation problems.
+DeDe can be applied to a wide range of resource allocation problems.
 Table 1 presents a summary of real-world resource allocation problems in the recent literature, mostly from OSDI, SOSP, NSDI, and SIGCOMM in the past four years.
-Next, we explain why Zeal is applicable.
+Next, we explain why DeDe is applicable.
   • Variables.
-Inheriting from ADMM, Zeal inherently supports real-number (float) allocation variables.
-Empirical evidence[53] and theoretical proofs [56, 38, 60] on ADMM also indicate that Zeal can effectively handle boolean and integer variables by projecting real-number solutions to desired data types over iterations.
+Inheriting from ADMM, DeDe inherently supports real-number (float) allocation variables.
+Empirical evidence[53] and theoretical proofs [56, 38, 60] on ADMM also indicate that DeDe can effectively handle boolean and integer variables by projecting real-number solutions to desired data types over iterations.
   • Objective.
-Zeal requires a max or sum of convex utility functions over per-resource or per-demand allocations.
+DeDe requires a max or sum of convex utility functions over per-resource or per-demand allocations.
 Most objectives used in systems exhibit convexity, such as the weighted allocation sum [44, 63], the logarithmic resource utility [3], and the quadratic allocation cost [64, 59].
   • Constraints.
-Zeal assumes the linearity of resource and demand constraints, which holds in all our surveyed problems, e.g., the sum of demands allocated to a resource must not exceed its capacity.
+DeDe assumes the linearity of resource and demand constraints, which holds in all our surveyed problems, e.g., the sum of demands allocated to a resource must not exceed its capacity.
 The constraints in these problems are per resource or per demand as well.
-Even under a constraint on multiple resources or demands, Zeal can be relaxed to groups of resources or demands.
-Zeal bundles these groups together into per-resource-group or per-demand-group allocations during decomposition.
-Unlike POP [43], Zeal works with any workload, not only “granular” ones.
+Even under a constraint on multiple resources or demands, DeDe can be relaxed to groups of resources or demands.
+DeDe bundles these groups together into per-resource-group or per-demand-group allocations during decomposition.
+Unlike POP [43], DeDe works with any workload, not only “granular” ones.
 In POP, the demands in each decomposed problem can access only a subset of resources, assuming each demand only requests a small fraction of interchangeable resources such that a resource subset suffices.
 This granular assumption can be unrealistic in practice, due to uneven demand distributions [28, 57] or limited resources that each demand is allowed to utilize [58, 33].
-In contrast, each decomposed problem in Zeal can still access all the resources or demands; thus, the granular assumption is no longer required, affirming Zeal’s generality to various workload types.
+In contrast, each decomposed problem in DeDe can still access all the resources or demands; thus, the granular assumption is no longer required, affirming DeDe’s generality to various workload types.
 4.2 Limitations
-While Zeal applies to a variety of resource allocation problems, it might not always be the most suitable solution for every allocation problem.
+While DeDe applies to a variety of resource allocation problems, it might not always be the most suitable solution for every allocation problem.
 Limited parallelism.
-Zeal achieves parallelism by decomposing the full problem into per-resource and per-demand allocations.
-Therefore, when we cannot fully decompose the problem, Zeal will experience limited parallelism, although this does not happen in our survey (Table 1).
+DeDe achieves parallelism by decomposing the full problem into per-resource and per-demand allocations.
+Therefore, when we cannot fully decompose the problem, DeDe will experience limited parallelism, although this does not happen in our survey (Table 1).
   • Non-separable objectives.
@@
 This also applies to demand constraints.
-Suboptimal solutions.  It is possible for Zeal not to attain an ideal solution when the allocation problem deviates from the standard ADMM formulation.
+Suboptimal solutions.  It is possible for DeDe not to attain an ideal solution when the allocation problem deviates from the standard ADMM formulation.
   • Non-convex problems.
 ADMM can solve non-convex optimization problems [53, 62, 68] empirically, providing theoretical guarantees only under specific assumptions[56, 38, 60].
-As a result, Zeal can output suboptimal solutions in certain non-convex cases.
+As a result, DeDe can output suboptimal solutions in certain non-convex cases.
   • Nonlinear constraints.
@@
 5 Case Studies
-In this section, we will present case studies on how Zeal can be applied to three resource allocation problems: traffic engineering, cluster scheduling, and load balancing, as also studied in POP [43].
+In this section, we will present case studies on how DeDe can be applied to three resource allocation problems: traffic engineering, cluster scheduling, and load balancing, as also studied in POP [43].
 5.1 Traffic Engineering
@@
 The problem shares the same variables and linear demand constraints as the max total flow problem.
-  • Objective.  The link utilization of a link $(u,v)$ is the ratio between the total flow allocated on that link and link capacity. Our objective is to minimize the maximum utilization of each link:
+  • Objective.  The utilization of a link $(u,v)$ is the ratio of the total flow assigned to the link to its capacity. Our objective is to minimize the max link utilization across the network:
     $\operatorname*{Minimize\leavevmode\nobreak\ }_{x}\max_{(u,v)\in E}\frac{\sum_{(s,t)}x_{(u,v)(s,t)}}{c_{u,v}}.$
-Using Zeal.
-The traffic engineering problem aligns the separable structure for Zeal: the objectives are the sum of allocation quality per demand or the maximum of allocation quality per link, respectively.
+Using DeDe.
+The traffic engineering problem aligns the separable structure for DeDe: the objectives are the sum of allocation quality per demand or the maximum of allocation quality per link, respectively.
 Linear constraints are applied to each demand and each link.
-Therefore, Zeal can decompose the problem into $|E|$ per-link allocations and $|V|^{2}$ per-demand allocations for parallelism.
+Therefore, DeDe can decompose the problem into $|E|$ per-link allocations and $|V|^{2}$ per-demand allocations for parallelism.
 5.2 Cluster Scheduling
@@
     $\operatorname*{Maximize\leavevmode\nobreak\ }_{x}\sum_{j}\log(\text{throughput}(j,x_{*j})).$
-Using Zeal.
-The cluster scheduling exhibits the separable structure required by Zeal: The objective function is a minimum of allocation quality per job and a sum of allocation quality per job, respectively.
+Using DeDe.
+The cluster scheduling exhibits the separable structure required by DeDe: The objective function is a minimum of allocation quality per job and a sum of allocation quality per job, respectively.
 Additionally, linear constraints are applied to each job and each computing resource.
-Thus, Zeal can decompose the problem into $n$ per-job allocations and $m$ per-demand allocations.
+Thus, DeDe can decompose the problem into $n$ per-job allocations and $m$ per-demand allocations.
 5.3 Load Balancing
@@
     $\sum_{i}x_{ij}=1,\quad\forall j.$
-Using Zeal.
-Given the separable structure, where the objective is the sum of allocation quality per data shard and linear constraints apply to each server, and each shard, Zeal can decompose the problem into $n$ per-server allocations and $m$ per-shard allocations.
-6 Implementation of Zeal
-We implement Zeal as a Python package. After publication, we will upload Zeal to PyPI for easy installation (pip install zeal).
-Listing 1 showcases a simple resource allocation problem solved using our Zeal package.
+Using DeDe.
+Given the separable structure, where the objective is the sum of allocation quality per data shard and linear constraints apply to each server, and each shard, DeDe can decompose the problem into $n$ per-server allocations and $m$ per-shard allocations.
+6 Implementation of DeDe
+We have implemented DeDe as a Python package and uploaded it to PyPI for easy installation: pip install dede.
+Listing 1 showcases a simple resource allocation problem solved using our DeDe package.
 In line 4, we initially create a $N\times M$ non-negative allocation variable x, where x[i,j] represents the allocation of resource i to demand j.
@@
 Finally, we construct the problem in line 16–17 and solve it in line 18.
-1import zeal
+1import dede
 2
-3# Create allocation variables.
-4x = zeal.Variable((N, M), nonneg=True)
+3# Create allocation variables
+4x = dede.Variable((N, M), nonneg=True)
 5
-6# Create the constraints.
+6# Create the constraints
 7resource_constrs = [
@@
 11
-12# Create an objective.
-13obj = zeal.Minimize(x.sum())
+12# Create an objective
+13obj = dede.Minimize(x.sum())
 14
-15# Construct and solve the problem.
-16prob = zeal.Problem(
+15# Construct and solve the problem
+16prob = dede.Problem(
 17 obj, resource_constrs, demand_constrs)
 18prob.solve(num_cpus=64)
-Listing 1: Resource allocation example with Zeal package.
-Zeal borrows the interface from cvxpy [17, 4].
-Zeal inherits most of cvxpy’s methods, e.g.
+Listing 1: Resource allocation example with DeDe package.
+DeDe borrows the interface from cvxpy [17, 4].
+DeDe inherits most of cvxpy’s methods, e.g.
 Variable(.) in line 4, Minimize(.) in line 13.
 One difference from cvxpy is that users must specify resource constraints and demands constraints separately (line 16–17) when initializing a problem.
-As a parallel solution, Zeal allows users to specify the number of CPU cores (line 18); if not specified, Zeal uses all available CPU cores.
-Behind Zeal’s interface, the resource allocation problem is solved in the following three steps.
+As a parallel solution, DeDe allows users to specify the number of CPU cores (line 18); if not specified, DeDe uses all available CPU cores.
+Behind DeDe’s interface, the resource allocation problem is solved in the following three steps.
   • Problem parsing.
@@
   1. 1.
-    How does Zeal compare with state-of-the-art approaches in allocation quality and computation time? (§7.1)
+    How does DeDe compare with state-of-the-art approaches in allocation quality and computation time? (§7.1)
   2. 2.
-    How do Zeal and state-of-the-art approaches react to changes in problem granularity, temporal fluctuation, spatial redistribution, and link failures? (§7.2)
+    How do DeDe and state-of-the-art approaches react to changes in problem granularity, temporal fluctuation, spatial redistribution, and link failures? (§7.2)
   3. 3.
-    What is the contribution of each design in Zeal? (§7.3)
+    What is the contribution of each design in DeDe? (§7.3)
 We focus on four approaches throughout our evaluation:
@@
 However, POP only simulates the parallel execution by solving subproblems sequentially, collecting the solving time of each, and calculating the parallel solving time mathematically.
-  • Zeal.  Zeal is our truly parallel implementation, with details described in Section 6.
-  • Zeal.
-To facilitate a fair comparison with POP, we create a variant of Zeal, denoted as Zeal, following POP-style simulation where subproblems are solved sequentially and a parallel solving time is calculated mathematically.
+  • DeDe.  DeDe is our truly parallel implementation, with details described in Section 6.
+  • DeDe.
+To facilitate a fair comparison with POP, we create a variant of DeDe, denoted as DeDe, following POP-style simulation where subproblems are solved sequentially and a parallel solving time is calculated mathematically.
 We also consider domain-specific approaches apart from these four approaches.
@@
 The computation time is measured on 64 CPU cores (2$\times$ Intel Xeon Gold 6142); an additional GPU (Nvidia Titan RTX) is made available for Teal.
-7.1 Zeal vs. the State of the Art
-In this section, we compare Zeal against state-of-the-art approaches on resource allocation problems from Section 5.
+7.1 DeDe vs. the State of the Art
+In this section, we compare DeDe against state-of-the-art approaches on resource allocation problems from Section 5.
 7.1.1 Traffic Engineering
@@
 Maximize total flow.
-Figure 3 compares Zeal to state-of-the-art methods in maximizing total flow.
-Zeal achieves 90.8% satisfied demand within 30 seconds and 92% within 60 seconds, demonstrating a superior trade-off between satisfied demand and computation time compared to Exact Sol, POP, and Pinning.
+Figure 3 compares DeDe to state-of-the-art methods in maximizing total flow.
+DeDe achieves 90.8% satisfied demand within 30 seconds and 92% within 60 seconds, demonstrating a superior trade-off between satisfied demand and computation time compared to Exact Sol, POP, and Pinning.
 While POP-4 also reaches 92% satisfied demand, its average computation time is significantly longer at 1658 seconds.
@@
 However, it relies on a highly customized machine-learning framework, demanding significant human effort for design and training.
-Zeal, representing the same algorithm under ideal conditions as POP, achieves a satisfied demand that is 8% higher than Zeal within 10 seconds in Figure 3 and consistently delivers superior allocation quality across the other figures.
-The reason is that Zeal executes each iteration more quickly, allowing it to complete more iterations within a given time, resulting in higher allocation quality.
-The difference in iteration time between Zeal and Zeal arises from three factors, as discussed in more detail in § 7.3: (1) Zeal measures the end-to-end time, including problem-solving, compilation, solution unpacking, and other associated tasks, while Zeal only takes the problem-solving time into account.
-(2) Zeal’s parallel implementation (§ 6) incurs overhead from cache contention, which slows down the problem-solving process.
-(3) In Zeal, each subproblem is statically pre-assigned to one of the processes, making it susceptible to stragglers.
-In contrast, Zeal assumes perfect dynamic scheduling.
+DeDe, representing the same algorithm under ideal conditions as POP, achieves a satisfied demand that is 8% higher than DeDe within 10 seconds in Figure 3 and consistently delivers superior allocation quality across the other figures.
+The reason is that DeDe executes each iteration more quickly, allowing it to complete more iterations within a given time, resulting in higher allocation quality.
+The difference in iteration time between DeDe and DeDe arises from three factors, as discussed in more detail in § 7.3: (1) DeDe measures the end-to-end time, including problem-solving, compilation, solution unpacking, and other associated tasks, while DeDe only takes the problem-solving time into account.
+(2) DeDe’s parallel implementation (§ 6) incurs overhead from cache contention, which slows down the problem-solving process.
+(3) In DeDe, each subproblem is statically pre-assigned to one of the processes, making it susceptible to stragglers.
+In contrast, DeDe assumes perfect dynamic scheduling.
 Figure 4: Comparison for the problem of minimizing max link utilization in traffic engineering.
 Min-max link utilization.
-Figure 4 compares Zeal with other approaches in minimizing maximum link utilization.
-Zeal achieves a low maximum link utilization of 1.67 within 10 seconds and 1.71 within 5 seconds.
+Figure 4 compares DeDe with other approaches in minimizing maximum link utilization.
+DeDe achieves a low maximum link utilization of 1.67 within 10 seconds and 1.71 within 5 seconds.
 Exact Sol reaches 1.63 with an average computation time of 35 seconds, and it is faster than itself in maximizing total flow because minimizing max link utilization is "easier" and requires fewer iterations.
-Pinning takes 42.5 seconds, making it slower than Zeal and slightly slower than Exact Sol, as it requires additional time to rebuild the problem when the top 10% of node pairs change.
+Pinning takes 42.5 seconds, making it slower than DeDe and slightly slower than Exact Sol, as it requires additional time to rebuild the problem when the top 10% of node pairs change.
 POP’s variants, POP-4, POP-16, and POP-64, achieve a maximum link utilization of 1.70, 1.77, and 1.95 with 9, 2, and 0.4 seconds on average, respectively.
-Although POP-16 is faster than Zeal, a fairer comparison between POP-16 and Zeal, both simulating parallelism, reveals that Zeal achieves a comparable computation time to POP-16.
+Although POP-16 is faster than DeDe, a fairer comparison between POP-16 and DeDe, both simulating parallelism, reveals that DeDe achieves a comparable computation time to POP-16.
 The domain-specific solution Teal takes only 0.3 seconds with massive parallelism from GPU with a maximum link utilization of 1.74.
@@
 Max-min allocation.
-Figure 5 compares Zeal against other approaches with the objective of max-min allocation.
-Zeal achieves a near-optimal max-min allocation of 0.97 within 5 seconds and 0.98 within 10 seconds, while Zeal reaches 0.98 within 2 seconds.
+Figure 5 compares DeDe against other approaches with the objective of max-min allocation.
+DeDe achieves a near-optimal max-min allocation of 0.97 within 5 seconds and 0.98 within 10 seconds, while DeDe reaches 0.98 within 2 seconds.
 Due to their sequential nature, Exact Sol and Gandiva encounter a slow computation time of 86 and 22 seconds.
 The greedy heuristic Gandiva even experiences a 0.56 low max-min allocation.
-POP-4 deliverers a comparable time-accuracy tradeoff as Zeal but is 3.4$\times$ slower than Zeal, which is simulated in the same way as POP did for a fair comparison.
+POP-4 deliverers a comparable time-accuracy tradeoff as DeDe but is 3.4$\times$ slower than DeDe, which is simulated in the same way as POP did for a fair comparison.
 A faster POP variant, POP-16, exhibits a lower max-min allocation of 0.94, as it becomes more difficult for jobs to assess the ideal GPU/CPU among the limited selection in smaller subproblems.
@@
 Proportional fairness.
-Figure 6 compares Zeal against other approaches in maximizing proportional fairness.
-As the optimization problem becomes more difficult with the convex logarithm function in the objective, both Zeal and Zeal achieve a greater speedup than POP by thoroughly breaking down the problem into smaller and easier pieces.
-Zeal and Zeal reach a proportional fairness score of over 1,000 in 300 seconds and 100 seconds, respectively.
+Figure 6 compares DeDe against other approaches in maximizing proportional fairness.
+As the optimization problem becomes more difficult with the convex logarithm function in the objective, both DeDe and DeDe achieve a greater speedup than POP by thoroughly breaking down the problem into smaller and easier pieces.
+DeDe and DeDe reach a proportional fairness score of over 1,000 in 300 seconds and 100 seconds, respectively.
 In contrast, POP-4 takes 1,825 seconds to reach the same score, while POP-16 and POP-64 remain below 900 despite their shorter computation times.
@@
 Minimize shard movements.
-Figure 8 compares Zeal with state-of-the-art approaches in minimizing shard movements.
-Zeal achieves an average of 24.1 shard movements in 15 seconds, slightly better than POP-4’s 25.4 shard movements within 16.9 seconds.
-In faster scenarios, Zeal achieves 35.1 shard movements in 2 seconds, while POP-4 achieves slightly fewer movements of 32.2 in 1.7 seconds.
-In a more fair comparison with POP, Zeal consistently costs 15–18% fewer shard movements than POP given a similar time limit.
+Figure 8 compares DeDe with state-of-the-art approaches in minimizing shard movements.
+DeDe achieves an average of 24.1 shard movements in 15 seconds, slightly better than POP-4’s 25.4 shard movements within 16.9 seconds.
+In faster scenarios, DeDe achieves 35.1 shard movements in 2 seconds, while POP-4 achieves slightly fewer movements of 32.2 in 1.7 seconds.
+In a more fair comparison with POP, DeDe consistently costs 15–18% fewer shard movements than POP given a similar time limit.
 Exact Sol achieves the minimum shard movements of 13.2 but is significantly slower, taking an average of 1,572 seconds due to the NP-hardness of solving the mixed-integer linear programming problem inherent to load balancing.
@@
 7.2 Robustness
-We evaluate the robustness of Zeal and other baselines against changes in problem granularity, temporal fluctuations, spatial redistribution, and link failures.
-In all these evaluations, Zeal is run for 30 seconds.
+We evaluate the robustness of DeDe and other baselines against changes in problem granularity, temporal fluctuations, spatial redistribution, and link failures.
+In all these evaluations, DeDe is run for 30 seconds.
 To mitigate the impact of the changing resources and demands, we normalize the satisfied demand by the optimal satisfied demand from Exact Sol.
@@
 To quantify resource interchangeability, we use the mean edge betweenness centrality [10, 11], which measures the average percentage of demands served by a given edge.
-As the mean edge betweenness centrality decreases, all approaches experience a decline in normalized satisfied demand, ranging from 0.8–2.1% for Zeal, 2.8–4.2% for Teal, 2.5–5.6% for Pinning, and 1–12.3% for POP.
-However, the maximum decrease in POP is 5.9$\times$ greater than Zeal because POP’s subproblems struggle to find alternative resources within their restricted subsets, whereas Zeal ’s subproblems retain access to the full set of resources.
+As the mean edge betweenness centrality decreases, all approaches experience a decline in normalized satisfied demand, ranging from 0.8–2.1% for DeDe, 2.8–4.2% for Teal, 2.5–5.6% for Pinning, and 1–12.3% for POP.
+However, the maximum decrease in POP is 5.9$\times$ greater than DeDe because POP’s subproblems struggle to find alternative resources within their restricted subsets, whereas DeDe’s subproblems retain access to the full set of resources.
 Teal exhibits a slight 1.3% increase in satisfied demand when the mean edge betweenness centrality is minimal.
@@
 Next, we randomly draw a sample from the newly created normal distribution and add it to each demand in every time slot, simulating temporal fluctuations.
-Figure 7(b) demonstrates that Zeal effectively manages temporal fluctuations, with changes in normalized satisfied demand limited to 1.3–3.5%.
+Figure 7(b) demonstrates that DeDe effectively manages temporal fluctuations, with changes in normalized satisfied demand limited to 1.3–3.5%.
 In contrast, Teal suffers a significant decrease of 12.8%, as it lacks prior exposure to the distribution of these fluctuating demands during training.
@@
 Specifically, we adjust the top 10% of demands, which initially account for 88.4% of the total volume, so that they take up 80%, 60%, 40%, and 20% of the total volume instead.
-Figure 7(c) shows that Zeal consistently fulfills the highest demand across all spatial distributions.
+Figure 7(c) shows that DeDe consistently fulfills the highest demand across all spatial distributions.
 Pinning’s performance decreases by 8.3%, as its strategy of focusing only on top demands relies on the presence of a heavy-tailed demand distribution.
@@
 7.3 Micro-benchmarks
-We conduct micro-benchmarks to evaluate the impact of Zeal’s design components, with all evaluations performed in maximizing total flow in traffic engineering.
+We conduct micro-benchmarks to evaluate the impact of DeDe’s design components, with all evaluations performed in maximizing total flow in traffic engineering.
 Speedup when varying CPU cores.
-Figure 9(a) evaluates the speedup of Zeal, Zeal, and Exact Sol with 1, 4, 16, and 64 CPU cores.
-Zeal achieves 61.7$\times$ speedup under 64 cores, where the slight gap occurs as we cannot exactly divide the total computation time evenly across 64 cores.
-Zeal also achieves an almost linear speedup when the number of CPU cores grows from 1 to 16.
-However, this linearity diminishes with 64 cores, where Zeal only achieves 18.2$\times$ speedup.
+Figure 9(a) evaluates the speedup of DeDe, DeDe, and Exact Sol with 1, 4, 16, and 64 CPU cores.
+DeDe achieves 61.7$\times$ speedup under 64 cores, where the slight gap occurs as we cannot exactly divide the total computation time evenly across 64 cores.
+DeDe also achieves an almost linear speedup when the number of CPU cores grows from 1 to 16.
+However, this linearity diminishes with 64 cores, where DeDe only achieves 18.2$\times$ speedup.
 This is because utilizing all 64 cores for parallelism incurs overhead from cache contention and stragglers, as discussed in Section 7.1.
-In contrast to the linear speedup of Zeal and Zeal, the speedup of Exact Sol is sublinear and marginal, achieving only 3.4$\times$ speeds with 64 CPU cores.
+In contrast to the linear speedup of DeDe and DeDe, the speedup of Exact Sol is sublinear and marginal, achieving only 3.4$\times$ speeds with 64 CPU cores.
 This is because of commercial solvers’ sequential nature, e.g., the simplex method [45] in linear programming requires thousands to millions of small steps to reach the optimal solution.
 Convergence rate.
-Both Teal and Zeal employ ADMM for parallel optimization, but their approaches differ.
+Both Teal and DeDe employ ADMM for parallel optimization, but their approaches differ.
 Teal’s ADMM adapts its computations based on matrix multiplication to leverage GPU calculation capabilities.
-Figure 9(b) illustrates the convergence rates of Teal and Zeal under various initialization.
-Zeal takes 30 seconds to surpass 90% satisfied demand.
-However, when initialized with Teal’s coarse neural network solution, Zeal achieves this within 20 seconds.
-Conversely, with naive initialization, Zeal requires more than 60 seconds.
+Figure 9(b) illustrates the convergence rates of Teal and DeDe under various initialization.
+DeDe takes 30 seconds to surpass 90% satisfied demand.
+However, when initialized with Teal’s coarse neural network solution, DeDe achieves this within 20 seconds.
+Conversely, with naive initialization, DeDe requires more than 60 seconds.
 Teal, benefiting from the massive parallelism of GPUs, reaches 90% satisfied demand within 10 seconds.
-However, its adapted ADMM limits its performance, capping its satisfied demand at 91% at 60 seconds compared to Zeal’s 92%.
+However, its adapted ADMM limits its performance, capping its satisfied demand at 91% at 60 seconds compared to DeDe’s 92%.
 Comparison to more optimization methods.
-Figure 9(c) investigates more constrained optimization methods other than Zeal’s ADMM method in maximizing the total flow in traffic engineering.
-Section 7.1 has covered the computation time of the optimal Exact sol, the suboptimal POP-64, and the near-optimal Zeal.
-As Zeal replaces constraints with additional terms in the objective function, we compare similar methods, including the penalty method and the augmented Lagrangian method.
-The augmented Lagrangian method is the same as in Zeal’s ADMM but without decoupling and decomposition, and thus it is over 3$\times$ slower than Zeal in reaching a satisfied demand of over 90%.
-The penalty method, which gradually increases penalty parameters from 0 to infinity to enforce the converted constraints in the objective function, converges slowly and is over 30$\times$ slower than Zeal in reaching a satisfied demand of over 90%.
+Figure 9(c) investigates more constrained optimization methods other than DeDe’s ADMM method in maximizing the total flow in traffic engineering.
+Section 7.1 has covered the computation time of the optimal Exact sol, the suboptimal POP-64, and the near-optimal DeDe.
+As DeDe replaces constraints with additional terms in the objective function, we compare similar methods, including the penalty method and the augmented Lagrangian method.
+The augmented Lagrangian method is the same as in DeDe’s ADMM but without decoupling and decomposition, and thus it is over 3$\times$ slower than DeDe in reaching a satisfied demand of over 90%.
+The penalty method, which gradually increases penalty parameters from 0 to infinity to enforce the converted constraints in the objective function, converges slowly and is over 30$\times$ slower than DeDe in reaching a satisfied demand of over 90%.
 8 Related Work
@@
 This key insight allows us to systematically decouple resource and demand constraints and decompose the original optimization problem into independent and parallel subproblems for allocating individual resources and demands.
-We implement this approach, Zeal, as a library to augment a widely-used, open-source solver (cvxpy), offering a push-button solution that accelerates large-scale resource allocation.
-Through extensive evaluation across diverse resource allocation tasks, we demonstrate that Zeal significantly improves scalability while closely approximating the quality of exact solutions.
-Appendix A Mathematical Representation of Zeal
+We implement this approach, DeDe, as a library to augment a widely-used, open-source solver (cvxpy), offering a push-button solution that accelerates large-scale resource allocation.
+Through extensive evaluation across diverse resource allocation tasks, we demonstrate that DeDe significantly improves scalability while closely approximating the quality of exact solutions.
+Appendix A Mathematical Representation of DeDe
 A resource allocation problem can be written as

### v3

@@
 Abstract
-Resource allocation is fundamental for cloud systems to ensure efficient resource sharing among tenants.
-However, the scale of such optimization problems has outgrown the capabilities of commercial solvers traditionally employed in production.
-To scale up resource allocation, prior approaches either tailor solutions to specific problems or rely on assumptions tied to particular workloads.
-In this work, we revisit real-world resource allocation problems and uncover a common underlying structure: a vast majority of these problems are inherently separable, i.e., they optimize the aggregate utility of individual resource and demand allocations, under separate constraints for each resource and each demand.
-Building on this insight, we develop DeDe, a general, scalable, and theoretically grounded framework for accelerating resource allocation through a decouple-and-decompose approach.
-DeDe systematically decouples entangled resource and demand constraints, thereby decomposing the overall optimization into alternating per-resource and per-demand allocations, which can then be solved efficiently and in parallel.
-We have implemented DeDe as a library extension to an open-source solver, maintaining a familiar user interface.
-Experimental results across three prominent resource allocation tasks—traffic engineering, cluster scheduling, and load balancing—demonstrate DeDe’s substantial speedups and robust allocation quality.
+Efficient resource allocation is essential in cloud systems to facilitate resource sharing among tenants.
+However, the growing scale of these optimization problems have outpaced commercial solvers commonly employed in production.
+To accelerate resource allocation, prior approaches either customize solutions for narrow domains or impose workload-specific assumptions.
+In this work, we revisit real-world resource allocation problems and uncover a common underlying structure: the vast majority of these problems are inherently separable, i.e., they optimize the aggregate utility of individual resource and demand allocations, under separate constraints for each resource and each demand.
+Building on this observation, we develop DeDe, a scalable and theoretically rooted optimization framework for large-scale resource allocation.
+At the core of DeDe is a decouple-and-decompose approach: it decouples entangled resource and demand constraints and thereby decomposes the overall optimization into alternating per-resource and per-demand subproblems that can be solved efficiently and in parallel.
+We have implemented and released DeDe as a Python package with a familiar modeling interface.
+Our experiments on three representative resource allocation tasks—cluster scheduling, traffic engineering, and load balancing—demonstrate that DeDe delivers significant speedups while generating higher-quality allocations.
+This work is accepted at USENIX OSDI 2025.
 1 Introduction
-Resource allocation is critical to the orchestration of multi-tenant systems, which are managed by today’s cloud providers on an unprecedented scale.
-It allows cloud services to be delivered with efficiency, reliability, and cost-effectiveness, by judiciously distributing various compute resources (e.g., CPUs and GPUs) and I/O resources (e.g., memory, storage, and network bandwidth) across a vast number of clients.
-Common applications of resource allocation include cluster scheduling [46, 27, 44, 61, 64], traffic engineering [24, 1, 67, 50, 31], and load balancing [16, 48, 52, 37].
-To meet varying user or application demands (jobs, traffic flows, and workloads), these systems extensively employ commercial optimization solvers like Gurobi [22] to compute the allocation of expensive cloud infrastructure (cluster resources, network links, and servers) worth millions or even billions of dollars.
-As cloud environments rapidly expand and diversify, the sheer scale of modern resource allocation problems has exceeded the capabilities of commercial solvers.
-These problems may involve millions of variables, which require solvers tens of minutes to hours to compute a solution when fast allocation is needed to maintain service quality.
-This “scalability crisis” in resource allocation has prompted a flurry of recent research [1, 43, 63, 47, 42] that trades off allocation quality for speed using heuristics, approximate algorithms, or machine learning.
-Nevertheless, these studies are limited to either a single domain (e.g., wide-area network traffic engineering [1, 63]) or a specific objective (e.g., max-min fairness [42]), requiring substantial manual effort to adapt to new scenarios.
-In contrast, POP [43] takes a general approach to accelerating assorted resource allocation problems.
-It partitions a large problem into a few (e.g., 16) subproblems, each with a subset of resources and demands, and solves them independently.
-However, this partition hinges on a “granular” assumption—each demand requests only a small fraction of mutually interchangeable resources.
-This assumption proves brittle with real workloads and degrades solution quality (§7.2).
-In this study, we re-examine real-world resource allocation problems through a different lens and aim to develop a general and scalable optimization framework without catering to specific problems or making workload-dependent assumptions.
-Our approach, DeDe, is theoretically grounded and enables fine-grained decomposition of the optimization problem, alternating between per-resource and per-demand allocations that can be solved in parallel (Figure 1).
+Resource allocation remains a critical and challenging problem today, especially as cloud providers operate multi-tenant systems on an unprecedented scale—these systems must ensure efficient and fair allocation of computing, storage, and networking resources across a large number of clients, in order to maintain the quality of cloud services.
+Common scenarios of resource allocation include cluster scheduling [47, 28, 45, 63, 66], traffic engineering [24, 1, 69, 51, 32], and load balancing [14, 49, 53, 38].
+To meet diverse and dynamic user demands (jobs, traffic flows, and queries), these systems extensively employ commercial optimization solvers like Gurobi [22] to solve linear programs (LP) or mixed-integer linear programs (MILP) that determine the allocation of valuable cloud resources (compute units, network links, and data nodes) worth millions to billions of dollars.
+However, as cloud environments continue to expand and diversify, the sheer scale of modern resource allocation problems has exceeded the capabilities of commercial solvers.
+These problems may involve millions of variables and take tens of minutes or even hours to solve, whereas fast allocation is necessary to meet service-level objectives.
+This “scalability crisis” has spurred a flurry of studies [1, 65, 48, 43] that trade off solution quality for computational speed, using heuristics, approximate algorithms, or machine learning.
+Nevertheless, these approaches are typically restricted to specific domains (e.g., wide-area network traffic engineering) or particular optimization objectives (e.g., max-min fairness), limiting their broader applicability.
+While POP [44] accelerates assorted allocation tasks, it hinges on a “granular” assumption that each demand requests only a small fraction of interchangeable resources.
+This premise proves brittle under realistic workloads, leading to degraded solution quality (§7).
+In this study, we revisit real-world resource allocation problems through a different lens, avoiding domain-specific designs or workload-dependent assumptions.
+We introduce DeDe  [^1] DeDe encapsulates our core technique “Decouple and Decompose”., a practical and scalable optimization framework for solving resource allocation problems.
+Rooted in optimization theory, DeDe allows the global allocation problem to be decomposed into alternating per-resource and per-demand subproblems (Figure 1), which can be solved more efficiently and in parallel without compromising solution quality.
 Figure 1: Overview of DeDe.
-This decomposition is made possible by an intrinsic, workload-agnostic problem structure that we uncovered after surveying dozens of real-world resource allocation problems in recent literature (Table 1).
-We discovered that a vast majority of these problems are inherently separable: they optimize the sum (or maximum) of utilities (measurement of solution quality) from allocating each individual resource and demand, while ensuring feasible allocation for individual resource and demand as well.111 We focus on individually separable resources and demands for simplicity and due to their widespread presence in real systems, but they can be relaxed to groups of resources and demands as we show in §4.1.
-In other words, the optimization objective comprises separate per-resource and/or per-demand utilities, and the constraints are also separate—they are either (1) a “resource constraint,” enforcing that the total demands mapped to each resource do not exceed the resource’s capacity, or (2) a “demand constraint,” enforcing that the total resources assigned to each demand do not surpass the demand’s needs.
-A formal definition of this separable structure is provided in §2, where we also show that real-world resource allocation is largely formulated with amenable mathematical properties (e.g., convex utilities, linear constraints) for tractability.
-[^2] Rare exceptions to this common structure are discussed in §4.2.
-Nonetheless, the uncovered “separability” in resource allocation problems does not directly allow us to actually separate them into independent subproblems that address individual resources and demands as depicted in Figure 1.
-This is because each resource not only appears in its own resource constraint but also participates in the demand constraints associated with every demand mapped to it; likewise for each demand.
-Ignoring the dependencies across subproblems and greedily solving each subproblem independently would fail to find feasible and high-quality allocations.
-Given the intertwined nature of resource and demand constraints, DeDe first decouples them as follows.
-Let $x$ denote the resource allocation matrix between $n$ resources and $m$ demands, where $x_{ij}$ represents the amount of resource $i$ assigned to demand $j$.
-The key mathematical trick in DeDe is to introduce a duplicate of the matrix $x$, denoted as $z$, along with an additional constraint $x=z$.
-Next, DeDe iteratively and alternately solves for $x$ subject only to resource constraints with $z$ fixed, and solves for $z$ subject only to demand constraints with $x$ fixed.
-In each iteration, DeDe further ensures that $x$ and $z$ do not deviate significantly from each other in order to satisfy $x=z$.
-This decoupling transformation lends itself to a classical constrained optimization framework called ADMM (Alternating Direction Method of Multipliers) [9].
-ADMM allows DeDe to embed each of the resource and demand constraints into the original objective as a penalty for constraint violation, thereby constructing augmented Lagrangians [23] as new optimization objectives.
-ADMM provides the theoretical foundation that enables alternating optimizations of these augmented Lagrangians with respect to $x$ and $z$, respectively.
-Next, DeDe further decomposes the above optimizations into per-resource and per-demand allocations.
-This is feasible owing to the separable structure we identified across resource allocation problems: the original objective is separable on each resource and demand variable, and so are the penalty terms converted by ADMM from constraints in the augmented Lagrangians.
-After decomposition, per-resource allocations involve solving $n$ subproblems, each with $m$ demand variables, and $m$ subproblems with $n$ resource variables for per-demand allocations.
-Consequently, the scale of optimization is greatly reduced from $n\times m$ variables to only $n$ or $m$ variables in each subproblem.
-These subproblems can be solved independently with massive parallelism using off-the-shelf solvers.
-Figure 2 illustrates the “decouple and decompose” workflow of DeDe.
-To demonstrate DeDe’s practicality, we have implemented it as a Python package, installable via pip install dede.
-DeDe is built on the widely used open-source solver cvxpy, and its interface is also designed to resemble cvxpy for ease of use (§6).
-Moreover, instead of simulating the parallel execution of subproblems on multiple CPU cores as previous studies did [1, 43], we have carefully optimized DeDe’s implementation using Ray [40], enabling efficient utilization of multiple cores without incurring significant overheads, such as cache contention.
-We evaluate DeDe in three prominent resource allocation problems—traffic engineering, cluster scheduling, and load balancing—and demonstrate its faster speed and higher allocation quality than the state of the art (§7).
-Compared with the best variant of POP [43] for each problem, DeDe achieves a 5.3% higher allocation quality and a 7.6$\times$ speedup in traffic engineering, 14% and 1.34$\times$ in cluster scheduling, and 5.3% and 1.13$\times$ in load balancing.
-Figure 2: Workflow of DeDe.
+Figure 2: DeDe decouples constraints using an auxiliary variable, thereby decomposing optimization into subproblems.
+This decomposition is enabled by an intrinsic, workload-agnostic problem structure that we identified after surveying dozens of real-world resource allocation problems in the literature (Table 1).
+We find that the vast majority of these problems are inherently separable: they optimize the total utility of individual resource and demand allocations, subject to separate constraints for each resource and demand to ensure feasibility.
+To be more concrete, we illustrate this problem structure using a simplified job scheduling example.
+[…]
+Formally, let $x_{ij}$ denote the fraction of the hour that job $j$ (demand) runs on GPU type $i$ (resource).
+Naturally, we have the demand constraints: $\forall j,\sum_{i}x_{ij}\leq 1$.
+On the other hand, for each GPU type, the total GPU hours requested by all jobs must not exceed the available GPU capacity.
+Thus, the resource constraints are: $\forall i,\sum_{j}\text{req}_{j}\cdot x_{ij}\leq\text{capacity}_{i}$.
+The objective is to maximize the total weighted average throughput, $\sum_{j}w_{j}\cdot(\sum_{i}\text{tput}_{ij}\cdot x_{ij})$, reflecting different job priorities.
+$\square$
+This example demonstrates the “separable” structure frequently encountered in practical resource allocation problems: resource constraints are specified for each resource and its associated demands, demand constraints are specified for each demand and its contributing resources, and the objective function sums the utility achieved by each demand (or resource).
+Despite this seemingly modular structure, however, such optimization problems remain computationally challenging for commercial solvers to solve at scale.
+The difficulty arises from each allocation variable $x_{ij}$ simultaneously affecting both resource and demand constraints, precluding straightforward decomposition and parallel optimization.
+To address this challenge, DeDe first decouples entangled resource and demand constraints through a mathematical reformulation.
+It introduces a copy of the allocation matrix $x$ as an auxiliary variable $z$, and rewrites the demand constraints (and per-demand terms in the objective) in terms of $z$, while adding a new constraint $x=z$.
+This transformation is mathematically equivalent, preserving the optimal solution.
+Moreover, it also lends itself to ADMM (Alternating Direction Method of Multipliers) [8], a constrained optimization method based on Lagrange multipliers [6].
+By incorporating constraints into the objective via multipliers and constructing the (augmented) Lagrangian (see §3.1), ADMM provides a theoretical framework that allows DeDe to alternatively and iteratively optimize the Lagrangian with respect to each block of variables—$x$ or $z$—while holding the other fixed.
+Building on the reformulation, DeDe further decomposes the optimization into subproblems along each resource and each demand.
+This step capitalizes on the identified separable structure, wherein the introduced Lagrangian can be partitioned across individual resources and demands (see §3.2).
+Consequently, instead of solving a monolithic optimization over all $n$ resources and $m$ demands, DeDe breaks the problem into $n$ per-resource subproblems, each with only $m$ demand variables, and $m$ per-demand subproblems, each with only $n$ resource variables.
+These smaller subproblems are amenable to independent and parallel optimization using off-the-shelf solvers.
+Figure 2 depicts the decouple-and-decompose workflow of DeDe.
+To demonstrate DeDe’s practical utility, we have implemented DeDe as a Python package, installable via pip install dede.
+Built on the popular open-source modeling language cvxpy [15, 3], DeDe provides familiar APIs (§6) aligned with cvxpy.
+Moreover, unlike prior efforts that simulate parallelism [1, 44], our implementation is optimized for real parallel execution using Ray [41], enabling efficient utilization of many CPU cores with minimal overhead.
+We evaluate DeDe on three representative resource allocation tasks—cluster scheduling, traffic engineering, and load balancing, demonstrating its faster speed and higher allocation quality relative to the state of the art (§7).
+Compared with the best variant of POP [44] in each domain, DeDe achieves a 7.3% improvement in allocation quality and a 3.1$\times$ speedup for cluster scheduling, 5.3% and 7.6$\times$ for traffic engineering, and 12.6% and 2.2$\times$ for load balancing.
 2 Real-World Resource Allocation Problems
-In modern multi-tenant systems, resources such as CPUs, memory, and bandwidth are shared among numerous users and applications.
-A central resource allocator often casts the distribution of resources as an optimization problem (e.g., linear programming) and repeatedly computes feasible solutions with off-the-shelf solvers to accommodate changing demands.
+In modern cloud systems, resources such as CPUs, memory, and network bandwidth are shared among numerous users and applications.
+A central resource allocator often casts the distribution of resources as an optimization problem (e.g., LP or MILP) and repeatedly computes feasible solutions using solvers to accommodate changing demands.
 However, these solvers are struggling to keep up with the increasing problem sizes that frequently occur in fast-growing systems.
-They may take several hours to solve an allocation problem that attempts to assign thousands of resources to thousands of demands—several orders of magnitude slower than desired (e.g., second-level SLAs, or service-level agreements).
-This challenge of scaling up resource allocation has spurred several recent works [1, 43, 63, 47, 42], but they either rely on domain-specific customizations or workload-dependent assumptions.
-In this paper, we instead seek to take a general (i.e., domain- and workload-agnostic) approach to scaling resource allocation.
-After surveying real-world resource allocation problems from recent literature (Table 1), we discover that nearly all these formulations can be transformed into a separable structure (§1), which we formally describe as follows.
+They may take up to several hours to solve an allocation problem that attempts to assign thousands of resources to thousands of demands.
+This is several orders of magnitude slower than desired (e.g., service-level objectives measured in seconds).
+This challenge of scaling up resource allocation has prompted a series of recent works [1, 44, 65, 48, 43], but they either rely on domain-specific customizations or workload-dependent assumptions.
+In this work, we instead seek to take a general (i.e., domain- and workload-agnostic) approach to scaling resource allocation.
+Separable problem structure.
+Upon surveying real-world resource allocation problems from the literature (Table 1), we find that nearly all these formulations can be transformed into a separable structure (§1) formally described as follows.
+(We discuss the generality and limitations in §4.)
   • Variables.
-An allocation matrix $x\in\mathcal{X}$ specifies how $n$ resources should be shared among $m$ demands, where $x_{ij}$ represents the quantity (or fraction) of resource $i$ assigned to demand $j$.
-Accordingly, the $i$th row of $x$, denoted as $x_{i*}=[x_{i1},x_{i2},\dots,x_{im}]$, constitutes the allocation vector of resource $i$, and the $j$th column $x_{*j}=[x_{1j},x_{2j},\dots,x_{nj}]$ encodes the allocation of demand $j$.
-The constraint set $\mathcal{X}$ is the Cartesian product of each $\mathcal{X}_{ij}$, which can be a set of real numbers or integers.
+An allocation matrix $x\in\mathcal{X}$ specifies the allocation of $n$ resources among $m$ demands, where $x_{ij}$ indicates the amount or fraction of resource $i$ assigned to demand $j$ (or vice versa).
+Accordingly, the $i$-th row of $x$, denoted as $x_{i*}=(x_{i1},x_{i2},\dots,x_{im})$, constitutes the allocation vector of resource $i$, while the $j$-th column $x_{*j}=(x_{1j},x_{2j},\dots,x_{nj})$ encodes the allocation of demand $j$.
+The constraint set $\mathcal{X}$ is the Cartesian product of the individual domains $\mathcal{X}_{ij}$, but it is often just non-negative real numbers or integers.
   • Objective.
-The objective specifies a metric to be optimized over an allocation matrix $x$, primarily consisting of a sum of utilities (or costs, to measure allocation quality) from allocating each resource ($x_{i*}$) and demand ($x_{*j}$):
+The objective defines a metric to be maximized or minimized over an allocation matrix $x$, consisting of a sum of utilities or costs that quantify the outcome from allocating each resource $x_{i*}$ or demand $x_{*j}$:
     $\sum_{i}f_{i}(x_{i*})+\sum_{j}g_{j}(x_{*j}),$ (1)
-    where $f_{i}(\cdot):\mathbb{R}^{m}\rightarrow\mathbb{R}$ and $g_{j}(\cdot):\mathbb{R}^{n}\rightarrow\mathbb{R}$ are utility (or cost) functions.
-They are typically designed to be convex to ensure practical tractability.
-Minimizing the maximum of utilities can also be converted into regular minimizing utilities (Equation 1) with additional constraints [5].
+    where $f_{i}(\cdot):\mathbb{R}^{m}\rightarrow\mathbb{R}$ and $g_{j}(\cdot):\mathbb{R}^{n}\rightarrow\mathbb{R}$ are utility or cost functions.
+These functions are typically convex (by nature or by choice), facilitating tractable optimization.
+A variety of objectives can be transformed into this form using standard optimization techniques.
+E.g., maximizing the minimum utility can be converted into maximizing an auxiliary “min utility” variable, with additional constraints ensuring that each utility is at least as large as this variable.
   • Constraints.
-To ensure that individual resources and demands are not over-provisioned (i.e., exceeding a resource’s capacity or a demand’s requirements), there is typically a constraint for each resource and each demand.
-We observe that these constraints are largely linear, presumably for tractability as well.
-For simplicity, we further convert inequality constraints into equality constraints using non-negative slack variables (details in Section 6).
-As such, the resource constraint on resource $i$ can be expressed as
-    $R_{i}x_{i*}=r_{i},$ (2)
-    and the demand constraint on demand $j$ can be written as
-    $D_{j}x_{*j}=d_{j}.$ (3)
-Despite the separable objective and separable constraints over each resource and each demand, large resource allocation problems (e.g., with millions of variables) still remain painfully slow to solve as they cannot actually be “separated” into smaller, concurrent per-resource and per-demand subproblems.
-The main obstacle is that resource and demand constraints are fundamentally coupled.
-Equations 2 and 3 show that each $x_{i,j}$ not only participates in the resource constraint (over $x_{i*}$) but also in the demand constraint (over $x_{*j}$).
-Next, we describe our solution to this challenge.
+To enforce feasible allocation, one or more constraints can be imposed on each resource and each demand (e.g., to prevent oversubscription of resources and overprovisioning of demands).
+For computational tractability, these constraints are typically linear.
+Following standard optimization practices, we transform inequality constraints into equivalent equality constraints using slack variables.
+Thus, without loss of generality, we express the $rc_{i}\geq 1$ resource constraints on each resource $i$ as
+    $R_{i}x_{i*}=r_{i},\quad\forall i,$ (2)
+    where $R_{i}\in\mathbb{R}^{rc_{i}\times m},r_{i}\in\mathbb{R}^{rc_{i}}$ are constraint parameters. Likewise, each demand $j$ is subject to $dc_{j}\geq 1$ linear constraints:
+    $D_{j}x_{*j}=d_{j},\quad\forall j,$ (3)
+    where $D_{j}\in\mathbb{R}^{dc_{j}\times n},d_{j}\in\mathbb{R}^{dc_{j}}$ are constraint parameters.
+Despite the separable objective and separate constraints over each resource and each demand, large resource allocation problems (e.g., with millions of variables) still remain painfully slow to solve as they cannot actually be separated into smaller, concurrent subproblems.
+The main obstacle is that resource and demand constraints are fundamentally entangled.
+As shown in Equations 2 and 3, each $x_{i,j}$ not only participates in the resource constraints (over $x_{i*}$) but also in the demand constraints (over $x_{*j}$).
+Figure 3: Toy scenario of the job scheduling example.
+[…]
+The left panel of Figure 3 shows the capacity of each GPU type ($\text{capacity}_{i}$) in GPU hours, alongside the throughput table ($\text{tput}_{ij}$).
+E.g., GPU type $C$ has a total capacity of 1.2 GPU hours, and assigning job 1 to it yields a throughput of 10 TPS.
+The right panel shows the optimal allocation matrix ($x_{ij}$), which satisfies all resource and demand constraints.
+E.g., the optimal solution schedules job 1 on (one unit of) GPU type $A$ for 0.8 hours, and on GPU type $C$ for 0.2 hours.
+The maximum average throughput achieved is 18.8 TPS.
+$\square$
 3 DeDe: Decouple and Decompose
-We present DeDe, a general, scalable, and theoretically rooted framework for optimizing real-world resource allocation problems.
-Leveraging the separable structure (§2) that underpins a vast majority of these problems, DeDe enables massive parallelism with near-optimal solutions through a decouple-and-decompose technique.
-A formal, mathematical representation of DeDe is included in Appendix A.
+We present DeDe, a practical and theoretically rooted optimization framework for solving large-scale, real-world resource allocation problems.
+By exploiting the inherently separable structure of these problems as described earlier, DeDe achieves massive parallelism while producing near-optimal solutions through a decouple-and-decompose approach.
 3.1 Decouple
-Resource and demand constraints being intertwined is the fundamental cause of complexity in resource allocation, as the individual allocation of any resource or demand will affect all other allocations.
-An optimal solution requires collectively allocating all resources to all demands, while a greedy strategy will lead to suboptimal or unfeasible solutions.
-Existing domain-agnostic solutions do not address this intrinsic problem nature of resource allocation.
-Production systems employ commercial solvers and rely on linear programming (LP) or mixed-integer linear programming (MILP) algorithms (e.g., simplex methods [45], barrier methods [30], or branch-and-bound methods [8]) to iteratively explore the feasible region—a high-dimensional search space dictated by coupled resource and demand constraints.
-The state of the art, POP [43], naively breaks intertwined constraints by partitioning resources and demands into a handful of subsets (e.g., 4–64), and then independently allocating each subset of resources to the demands within the same subset.
-Doing so ignores the dependencies across subsets and leads to suboptimal solutions in practice (§7.1).
-Meanwhile, it only enables limited parallelism because finer-grained decomposition would render the allocation unfeasible and unusable.
-Instead of disregarding the constraints across subproblems, we tackle the intertwining root cause directly.
-Specifically, we aim to decouple demand constraints from resource constraints, transforming the full allocation into alternating per-resource allocations—assigning a resource to all demands subject to the corresponding resource constraint, and per-demand allocations—mapping a demand to all resources are subject to the corresponding demand constraint.
-To begin with, we offload demand constraints to another variable different from $x$.
-We create a copy of the allocation matrix $x$, denoted as $z$, and replace the variables in each demand constraint (Equation 3) with their counterparts in $z$:
-$D_{j}z_{*j}=d_{j}.$ (4)
-We also substitute $z$ for the demand-related utilities in the objective function (Equation 1):
-$\sum_{i}f_{i}(x_{i*})+\sum_{j}g_{j}(z_{*j}).$ (5)
-Then, representing the same allocation plan, $x$ and $z$ should be equal. Thus, we introduce a new constraint:
-$x=z.$ (6)
-Consequently, all resource constraints remain on $x$, but all demand constraints have been moved to $z$.
-The decoupling of constraints clearly leads to a formulation equivalent to the original optimization with the new constraint $x=z$.
-However, the new constraint still has $x$ and $z$ entangled, so the introduction of $z$ does not immediately enable decomposition.
-In fact, jointly optimizing $x$ and $z$ (e.g., using penalty methods [6] or the augmented Lagrangian method [23]) will forfeit the benefits of decoupling, as we will show in §7.3.
-On the bright side, the transformed formulation becomes compatible with a half-century-old constrained optimization framework called ADMM (Alternating Direction Method of Multipliers) [9], which provides theoretically guaranteed solutions when solving for two sets (“blocks”) of variables ($x$ and $z$).
-At a high level, the theory behind ADMM enables alternating optimization of the following subproblems without making compromises on the solution quality: minimizing the updated objective (Equation 5) with respect to
-  • $x$ (with $z$ fixed), subject to resource constraints and $x=z$,
-  • $z$ (with $x$ fixed), subject to demand constraints and $x=z$.
-We elaborate in §3.2 that both optimization subproblems are amenable to decomposition, reducing problem sizes and enabling massive parallelism.
-Next, we first describe how ADMM solves them (even without decomposition).
-  • Constraint transformation.
-ADMM is a variant of the augmented Lagrangian method [23].
-Taking the first subproblem above (for $x$) as an example: ADMM converts every constraint, i.e., resource constraints and $x=z$, into penalty terms and appends them to the objective to minimize (Equation 5).
-A conventional penalty method will add a single penalty term $\sum_{i}\mu\cdot c_{i}^{2}$ to the objective, where $c_{i}$ is a measure of the constraint violation (comprising $|R_{i}x_{i*}-r_{i}|$ and $|x_{i}-z_{i}|$) for resource $i$, and $\mu$ is a penalty coefficient that increases to infinity over iterations to ensure asymptotic convergence to the optimal solution.
-However, taking $\mu$ to infinity requires careful manual tuning, or otherwise incurs ill-conditioning [2].
-The augmented Lagrangian method has a similar rationale to penalty methods, except that it introduces yet another penalty term $\lambda_{i}\cdot c_{i}$ to mimic a Lagrange multiplier.
-The update of $\lambda_{i}$ is automated as $\lambda_{i}\leftarrow\lambda_{i}+\mu\cdot c_{i}$ with theoretical guarantees, where $\mu$ can stay as a small constant while avoiding ill-conditioning.
-  • Partial optimization.
-The new objectives with additional penalty terms converted from constraint violations are also called augmented Lagrangians.
-ADMM alternately minimizes these augmented Lagrangians, solving for $x$ with $z$ fixed, and solving for $z$ with $x$ fixed.
-This is known as “two-block partial optimization” and guarantees the convergence to optimal solutions for convex problems [21, 20].
-For non-convex problems (including those with integer variables), practical evidence suggests that ADMM is still effective at solving (some of) them as well [53, 62, 68].
-Theoretical proofs are available under specific assumptions [56, 38, 60].
+The interdependence between resource and demand constraints is a core complexity in resource allocation.
+Achieving an optimal solution requires jointly allocating all resources to all demands; otherwise, local, greedy strategies likely result in suboptimal or infeasible solutions.
+Existing solutions do not address this intertwined structure of constraints in resource allocation.
+In practice, production systems tend to simply rely on commercial solvers and employ linear programming (LP) or mixed-integer linear programming (MILP) algorithms (e.g., simplex methods [46], barrier methods [31], or branch-and-bound methods [7]) to navigate a high-dimensional search space, which is induced by the coupling of resource and demand constraints.
+A recent work, POP [44], naively breaks entangled constraints by partitioning resources and demands into a handful of subsets (e.g., 4–64), and then independently allocates a subset of resources to a subset of demands.
+Doing so ignores the dependencies across subsets and leads to suboptimal solutions in practice (§7).
+Meanwhile, it only enables limited parallelism because finer-grained decomposition would render the allocation infeasible and unusable.
+Instead of disregarding the constraints across subproblems, we tackle the root cause of constraint entanglement.
+Specifically, without sacrificing solution quality, our high-level goal is to decouple demand constraints from resource constraints, transforming the global allocation into alternating per-resource and per-demand subproblems.
+Reformulation.
+Consider the minimization of the objective in Equation 1 (the signs of $f_{i}$ and $g_{j}$ can be simply reversed for maximization problems), subject to resource constraints in Equation 2 and demand constraints in Equation 3.
+First, we introduce an auxiliary variable $z$ as a duplicate of the allocation matrix $x$, and rewrite all demand constraints and all demand-related terms in the objective using $z$.
+Meanwhile, we add a new constraint $x=z$.
+This transformation clearly leads to an equivalent optimization problem:
+$\min_{x\in\mathcal{X}}\quad$ $\sum_{i}f_{i}(x_{i*})+\sum_{j}g_{j}(z_{*j}),$
+$s.t.\quad$ $R_{i}x_{i*}=r_{i},\quad\forall i,$ (4)
+$D_{j}z_{*j}=d_{j},\quad\forall j,$
+$x-z=0.$
+In this formulation, all resource constraints remain on $x$, while all demand constraints (and related utilities) are moved to $z$.
+Although introducing $z$ appears to separate the problem structurally, the (necessary) constraint $x=z$ still tightly couples the variables and thus prevents straightforward decomposition.
+In fact, jointly optimizing $x$ and $z$ (e.g., using penalty methods [4] or augmented Lagrangian [23]) will forfeit the benefits of this reformulation, as we will show in §7.3.
+ADMM background.
+Nevertheless, this reformulation lends itself to ADMM (Alternating Direction Method of Multipliers) [8], a classical method for solving constrained optimization problems of the form: minimize $f(x)+g(z)$, subject to $Ax+Bz=c$.
+Note that Equation 3.1 conforms to this form.
+ADMM is based on augmented Lagrangian [23], a method that converts constrained optimization into a sequence of unconstrained optimization problems using Lagrange multipliers [6] (a well-known technique in optimization).
+These unconstrained problems are defined over the Lagrangian function, constructed from the objective and constraints: $L_{\rho}(x,z,y)=f(x)+g(z)+y^{\mathsf{T}}(Ax+Bz-c)+(\rho/2)\lVert Ax+Bz-c\rVert^{2}_{2}$, where $y$ is (an estimate of) the Lagrange multiplier, and the last quadratic term with parameter $\rho$ serves as a penalty [4] for improving convexity and stability.
+The theory of ADMM enables partial, alternating optimization of the Lagrangian with respect to $x$ and $z$—iteratively solving for $x$ with $z$ fixed, and solving for $z$ with $x$ fixed.
+This “two-block partial optimization” guarantees convergence to an optimal solution for convex problems [21, 20].
+In non-convex settings (including those involving integers), empirical results indicate that ADMM can still effectively solve many of these problems [54, 64, 70].
+Theoretical convergence results are established under certain conditions [57, 39, 61].
+By introducing the auxiliary variable $u=(1/\rho)y$, the augmented Lagrangian can be further rewritten as: $L_{\rho}(x,z,u)=f(x)+g(z)+(\rho/2)\lVert Ax+Bz-c+u\rVert^{2}_{2}-(\rho/2)\lVert u\rVert^{2}_{2}$, leading to a more convenient scaled form of ADMM.
+Next, we present the ADMM iterates for updating $x$, $y$, and $u$ directly in the context of resource allocation problems.
+Applying ADMM.  DeDe applies the scaled form of ADMM to solve the reformulated optimization problem (Equation 3.1). First, it constructs the augmented Lagrangian:
+$\begin{split}&\mathcal{L}_{\rho}(x,z,\alpha,\beta,\lambda)=\sum_{i}f_{i}(x_{i*})+\sum_{j}g_{j}(z_{*j})\\ &+\frac{\rho}{2}\Bigl{(}\sum_{i}\lVert R_{i}x_{i*}-r_{i}+\alpha_{i}\rVert_{2}^{2}+\sum_{j}\lVert D_{j}z_{*j}-d_{j}+\beta_{j}\rVert_{2}^{2}\\ &+\lVert x-z+\lambda\rVert_{F}^{2}\Bigr{)}-\frac{\rho}{2}(\lVert\alpha\rVert_{2}^{2}+\lVert\beta\rVert_{2}^{2}+\lVert\lambda\rVert_{F}^{2}),\end{split}$ (5)
+where $\alpha_{i}\in\mathbb{R}^{rc_{i}}$, $\beta_{j}\in\mathbb{R}^{dc_{j}}$, $\lambda\in\mathbb{R}^{n\times m}$ are the introduced auxiliary variables.
+$\lVert\cdot\rVert_{2}$ still denotes the $\ell^{2}$-norm of a vector, while $\lVert\cdot\rVert_{F}$ is the Frobenius norm of a matrix, i.e., the $\ell^{2}$-norm of the matrix when flattened into a vector.
+The ADMM iterates are given by:
+$x^{(k+1)}$ $:=\operatorname*{arg\,min}_{x\in\mathcal{X}}\,\mathcal{L}_{\rho}(x,z^{(k)},\alpha^{(k)},\beta^{(k)},\lambda^{(k)})$ (6)
+$z^{(k+1)}$ $:=\operatorname*{arg\,min}_{z}\,\mathcal{L}_{\rho}(x^{(k+1)},z,\alpha^{(k)},\beta^{(k)},\lambda^{(k)})$ (7)
+$\alpha_{i}^{(k+1)}$ $:=\alpha_{i}^{(k)}+R_{i}x^{(k+1)}_{i*}-r_{i},\forall i$
+$\beta_{j}^{(k+1)}$ $:=\beta_{j}^{(k)}+D_{j}z^{(k+1)}_{*j}-d_{j},\forall j$
+$\lambda^{(k+1)}$ $:=\lambda^{(k)}+x^{(k+1)}-z^{(k+1)}$
+We will show in §3.2 that both the optimization subproblems in Equations 6 and 7 are amenable to decomposition owing to the previously identified separable problem structure.
 3.2 Decompose
-To this point, the original resource allocation problem has been transformed into alternating minimization of two augmented Lagrangians (original objective plus penalty terms) with respect to $x$ and $z$, respectively.
-Next, we take the $x$-minimization step as an example to describe how it can be decomposed ($z$-minimization is similar).
-As $x$-minimization focuses on the variable $x$ only, we can ignore terms related to $z$ in the augmented Lagrangian, such as the penalty terms from demand constraints.
-The resulting objective consists of three resource-related terms:
-  • sum of per-resource allocation cost $f_{i}(x_{i*})$ from Equation 5;
-  • sum of penalties for violating each resource constraint $R_{i}x_{i*}=r_{i}$ from Equation 2;
-  • sum of penalties for violating $x=z$ from Equation 6.
-The first two items are a sum of functions on per-resource allocation $x_{i*}$, while the third can also be rewritten as a sum of penalties for violating each $x_{i*}=z_{i*}$ (equivalent to $x=z$).
-After rearranging, the optimization can be expressed as:
-$\min_{x\in\mathcal{X}}\sum_{i}F_{i}(x_{i*}),$
-where $F_{i}(x_{i*})$ consists of allocation cost, resource constraint penalty, and penalty for violating $x_{i*}=z_{i*}$.
-As each variable $x_{i*}$ is mutually exclusive and independent, the above problem is equal to optimizing per-resource allocations in parallel:
-$\min_{x_{i*}\in\mathcal{X}_{i*}}F_{i}(x_{i*}),\quad\forall i,$
-thereby decomposing the allocation of $x$ with $n\times m$ variables into $n$ per-resource allocations of $x_{i*}$, each with $m$ variables.
-To solve them, DeDe uses an off-the-shelf solver (cvxpy in our implementation) to find (exact or approximate) solutions while making liberal use of all available CPU cores.
+Up to this point, we have reformulated the original resource allocation problem into alternating minimization steps of the augmented Lagrangian with respect to $x$ and $z$.
+Next, we take the $x$-minimization step as an example to describe how the optimization is decomposed ($z$-minimization is similar).
 Variables Objective
@@
 PODP[5], RAS[47], Skyplane[27], Oort[35], TACCL[50], Shard Manager[38], Zeta[67], CASCARA[51], Sia[28], POP[44]
-NetHint [14], Gavel [44], Teal [63], OneWAN [32], BLASTSHIELD [31], NCFlow [1], Cerebro [66], DOTE [47], Soroush  [42], POP [43]
+NetHint[12], Gavel[45], Teal[65], ONEWAN[33], BLASTSHIELD[32], NCFlow[1], Cerebro[68], DOTE[48], POP[44]
 PCF[30], Electricity Pricing[60], POP[44]
 Table 1: Real-world resource allocation problems in recent literature.
+Because the $x$-minimization step solves for $x$ only, we can ignore all terms in the augmented Lagrangian without $x$ (such as demand constraints). Thus, we rewrite Equation 6 as
+$x^{(k+1)}=$ $\operatorname*{arg\,min}_{x\in\mathcal{X}}\,\mathcal{L}_{\rho}(x,z^{(k)},\alpha^{(k)},\beta^{(k)},\lambda^{(k)})$ (6)
+$=$ $\operatorname*{arg\,min}_{x\in\mathcal{X}}\,\sum_{i}f_{i}(x_{i*})+\frac{\rho}{2}\sum_{i}\big{\|}R_{i}x_{i*}-r_{i}+\alpha_{i}^{(k)}\big{\|}_{2}^{2}$
+$+\frac{\rho}{2}\big{\|}x-z^{(k)}+\lambda^{(k)}\big{\|}_{F}^{2}$
+$=$ $\operatorname*{arg\,min}_{x\in\mathcal{X}}\,\sum_{i}\Bigl{(}f_{i}(x_{i*})+\frac{\rho}{2}\big{\|}R_{i}x_{i*}-r_{i}+\alpha_{i}^{(k)}\big{\|}_{2}^{2}$
+$+\frac{\rho}{2}\big{\|}x_{i*}-z_{i*}^{(k)}+\lambda_{i*}^{(k)}\big{\|}_{2}^{2}\Bigr{)}.$
+Given that the new objective is a sum of disjoint terms across index $i$, its minimization can be decomposed into independent, per-resource subproblems over each $x_{i*}$:
+$\begin{split}x_{i*}^{(k+1)}=&\operatorname*{arg\,min}_{x_{i*}\,\in\mathcal{X}_{i*}}\,\Bigl{(}f_{i}(x_{i*})+\frac{\rho}{2}\big{\|}R_{i}x_{i*}-r_{i}+\alpha_{i}^{(k)}\big{\|}_{2}^{2}\\ &+\frac{\rho}{2}\big{\|}x_{i*}-z_{i*}^{(k)}+\lambda_{i*}^{(k)}\big{\|}_{2}^{2}\Bigr{)},\quad\forall i.\end{split}$ (8)
+This decomposition is made possible by the separable structure: (1) the original objective is separable, with a utility $f_{i}(x_{i*})$ defined for allocating each resource $i$, and (2) the resource constraints are likewise separable, as each constraint $R_{i}x_{i*}=r_{i}$ pertains only to the corresponding resource $i$.
+As a result, the original optimization problem over $n\times m$ variables is broken down into $n$ independent subproblems, each involving only $m$ variables.
+DeDe solves each subproblem using off-the-shelf solvers, which return either exact or approximate solutions depending on the nature of $f$ (convex or not), whether $\mathcal{X}$ contains integers, etc.
+Since this decomposition allows up to $n$ parallel computations—often numbering in the thousands in large-scale settings—DeDe is capable of delivering substantial speedups assuming sufficient CPUs.
+For linear programs, the original problem has a worst-case time complexity of approximately $O((n\cdot m)^{2.373})$ [13, 44].
+After decomposition, DeDe processes $n$ subproblems, each with a complexity of $O(m^{2.373})$, reducing the overall complexity to $O(n\cdot m^{2.373})$.
+Nonetheless, this theoretical improvement assumes that ADMM converges within a constant number of iterations.
+In practice, the number of iterations depends on the choice of the penalty parameter $\rho$ (which acts like a “learning rate”) and other factors that influence convergence.
+Similarly, the $z$-minimization step decomposes into independent, per-demand allocation subproblems over each $z_{*j}$:
+$\begin{split}z_{*j}^{(k+1)}=&\operatorname*{arg\,min}_{z_{*j}}\,\Bigl{(}g_{j}(z_{*j})+\frac{\rho}{2}\big{\|}D_{j}z_{*j}-d_{j}+\beta_{j}^{(k)}\big{\|}_{2}^{2}\\ &+\frac{\rho}{2}\big{\|}x^{(k+1)}_{*j}-z_{*j}+\lambda_{*j}^{(k)}\big{\|}_{2}^{2}\Bigr{)},\quad\forall j.\end{split}$ (9)
 4 Generality and Limitations
-In this section, we start with the generality of DeDe, and then discuss its limitations, including scenarios with limited parallelism and chances of suboptimal solutions.
+In this section, we discuss the generality of DeDe (conditions under which it applies) and its limitations (scenarios that may result in limited parallelism or suboptimal solutions).
 4.1 Generality
-DeDe can be applied to a wide range of resource allocation problems.
-Table 1 presents a summary of real-world resource allocation problems in the recent literature, mostly from OSDI, SOSP, NSDI, and SIGCOMM in the past four years.
-Next, we explain why DeDe is applicable.
+DeDe is applicable to a wide range of real-world resource allocation problems.
+Table 1 summarizes these problems that appear in recent literature such as OSDI, SOSP, NSDI, and SIGCOMM.
+Below, we explain why DeDe is applicable.
   • Variables.
-Inheriting from ADMM, DeDe inherently supports real-number (float) allocation variables.
-Empirical evidence[53] and theoretical proofs [56, 38, 60] on ADMM also indicate that DeDe can effectively handle boolean and integer variables by projecting real-number solutions to desired data types over iterations.
+As an ADMM-based framework, DeDe natively supports continuous (real-valued) allocation variables.
+Empirical evidence[54] and theoretical proofs [57, 39, 61] on ADMM further indicate that DeDe should effectively handle boolean and integer variables by projecting real-valued solutions onto the appropriate domains during iterations.
   • Objective.
-DeDe requires a max or sum of convex utility functions over per-resource or per-demand allocations.
-Most objectives used in systems exhibit convexity, such as the weighted allocation sum [44, 63], the logarithmic resource utility [3], and the quadratic allocation cost [64, 59].
+Most objectives in real systems are defined as the sum, max, or min of utility functions over the individual allocation of resources and demands.
+They also exhibit convexity.
+Examples include weighted sums [45, 65], logarithmic utilities [2], and quadratic costs [66, 60].
   • Constraints.
-DeDe assumes the linearity of resource and demand constraints, which holds in all our surveyed problems, e.g., the sum of demands allocated to a resource must not exceed its capacity.
-The constraints in these problems are per resource or per demand as well.
-Even under a constraint on multiple resources or demands, DeDe can be relaxed to groups of resources or demands.
-DeDe bundles these groups together into per-resource-group or per-demand-group allocations during decomposition.
+DeDe assumes linear constraints on resources and demands—a condition that holds across all the surveyed use cases.
+These constraints are commonly expressed at the per-resource or per-demand level.
+Even for constraints that do span multiple resources or demands, DeDe can turn them into disjoint constraint groups and apply decomposition at the group level (at the cost of reduced parallelism).
 Unlike POP [44], DeDe works with any workload, not only “granular” ones.
-In POP, the demands in each decomposed problem can access only a subset of resources, assuming each demand only requests a small fraction of interchangeable resources such that a resource subset suffices.
-This granular assumption can be unrealistic in practice, due to uneven demand distributions [28, 57] or limited resources that each demand is allowed to utilize [58, 33].
-In contrast, each decomposed problem in DeDe can still access all the resources or demands; thus, the granular assumption is no longer required, affirming DeDe’s generality to various workload types.
+In POP, the demands in each subproblem can only access a subset of resources, assuming each demand only requests a small fraction of interchangeable resources such that a resource subset suffices.
+This granular assumption can be unrealistic in practice, due to uneven demand distributions [29, 58] or specific resources that each demand is allowed to utilize [59, 34].
+In contrast, each subproblem in DeDe can still access all the resources or demands, resulting in generalization to various workload distributions.
 4.2 Limitations
-While DeDe applies to a variety of resource allocation problems, it might not always be the most suitable solution for every allocation problem.
+While DeDe applies to a variety of resource allocation tasks, it might not always be the optimal approach for all scenarios.
 Limited parallelism.
-DeDe achieves parallelism by decomposing the full problem into per-resource and per-demand allocations.
-Therefore, when we cannot fully decompose the problem, DeDe will experience limited parallelism, although this does not happen in our survey (Table 1).
+DeDe’s parallelism relies on its ability to decompose the problem into independent subproblems.
+When such decomposition is not fully achievable, the degree of parallelism diminishes.
+Although this limitation does not manifest in the set of real-world problems we surveyed, we come up with some examples below.
   • Non-separable objectives.
-When the objective function is not a sum or max over allocation qualities per resource or per demand, we must bundle per-resource or per-demand allocations whose utilities stick together during decomposition.
-As a result, the partially decomposed problem will achieve limited parallelism.
+An objective is non-separable when utility depends on the interaction between different resources or demands, e.g., if a job attains higher throughput when it has exclusive access to a GPU compared with when sharing it with other jobs.
   • Non-separable constraints.
-When a resource constraint restricts the allocation across multiple resources, we must bundle the allocations for these resources during decomposition, which leads to limited parallelism.
-This also applies to demand constraints.
-Suboptimal solutions.  It is possible for DeDe not to attain an ideal solution when the allocation problem deviates from the standard ADMM formulation.
+A constraint is non-separable when it spans multiple resources and multiple demands simultaneously.
+For example, this occurs when the jobs submitted by each user are collectively limited by a user-specific GPU hour quota.
+To restore decomposability, DeDe has to treat all jobs (the original demands) from the same user as a single aggregated demand.
+However, this aggregation reduces the granularity of parallelism, from the job level to the user level.
+Suboptimal solutions.  DeDe may fail to reach the optimal solution when the allocation problem deviates from the assumptions underlying standard ADMM.
   • Non-convex problems.
-ADMM can solve non-convex optimization problems [53, 62, 68] empirically, providing theoretical guarantees only under specific assumptions[56, 38, 60].
-As a result, DeDe can output suboptimal solutions in certain non-convex cases.
+Although ADMM has demonstrated empirical success on various non-convex problems [54, 64, 70], its theoretical convergence is guaranteed only under specific conditions [57, 39, 61].
+Consequently, DeDe may converge to suboptimal solutions in non-convex settings.
   • Nonlinear constraints.
-When a constraint is nonlinear, we replace it by adding an indicator function to the objective.
-The indicator function is infinity when the constraint is violated or otherwise zero.
-However, this may alter the convexity of the objective function, resulting in the above non-convex problems.
+When a constraint is nonlinear, DeDe has to reformulate it with an indicator function in the augmented Lagrangian.
+The indicator function returns zero if the constraint is satisfied and infinity otherwise.
+However, this approach destroys the convexity of the objective, resulting in the aforementioned non-convex problems.
   • Higher allocation dimension.
-The allocation of resources to demands is encoded in a two-dimensional matrix $x$.
-Correspondingly, we separate the problem into two ADMM blocks for resources and demands.
-Introducing more dimensions, such as a temporal dimension [36], leads to more ADMM blocks and risks convergence failure [13].
+DeDe assumes a two-dimensional allocation matrix that maps resources to demands, leading to a two-block ADMM formulation.
+Introducing additional dimensions such as a temporal dimension [37] will increase the number of variable blocks in ADMM and thus might lead to convergence issues [11].
 5 Case Studies
-In this section, we will present case studies on how DeDe can be applied to three resource allocation problems: traffic engineering, cluster scheduling, and load balancing, as also studied in POP [43].
+In this section, we present case studies of DeDe applied to three representative resource allocation problems [44]: cluster scheduling, traffic engineering, and load balancing.
+5.1 Cluster Scheduling
+Cluster scheduling problems [63, 45, 44, 43] involve allocating computing jobs to different resource types in heterogeneous clusters.
+The job scheduling example introduced earlier is a simplified version of this problem.
+Max-min allocation.  This variant aims to distribute the heterogeneous cluster resources fairly among jobs such that the minimum throughput is maximized.
+  • Variables.
+Jobs can be time-sliced across available resource types [45, 44].
+Thus, the allocation plan is a matrix $x\in[0,1]^{n\times m}$, where $x_{ij}$ is the fraction of time (optimization interval) that each job $j$ spends on each resource type $i$.
+  • Objective.
+Given an allocation matrix $x$, we adopt a normalized effective throughput of job $j$ from POP [42], which is a linear function of $x_{*j}$.
+The objective is to maximize the minimum normalized effective throughput:
+    $\operatorname*{Maximize\leavevmode\nobreak\ }_{x}\min_{j}\leavevmode\nobreak\ \text{throughput}(j,x_{*j}).$
+  • Resource constraints.  For each computing resource type $i$, the total amount of resources requested by all jobs cannot exceed the corresponding capacity:
+    $\sum_{j}x_{ij}z_{j}\leq\text{capacity}_{i},\quad\forall i,$
+    where $z_{j}$ is the amount of resources requested by job $j$.
+  • Demand constraints.  For each job $j$, the sum of all fractions of time must not exceed 1 naturally:
+    $\sum_{i}x_{ij}\leq 1,\quad\forall j.$
+Proportional fairness.
+The second variant of cluster scheduling maximizes the overall resource utilization while ensuring minimum service for each job.
+This variant shares the same variables and constraints as above, except for the objective:
+  • Objective.  Proportional fairness aims to maximize the sum of log utilities of each job:
+    $\operatorname*{Maximize\leavevmode\nobreak\ }_{x}\sum_{j}\log(\text{throughput}(j,x_{*j})).$
+Applying DeDe.
+Both variants of cluster scheduling exhibit the separable structure required by DeDe.
+Their objectives to maximize are a sum or min of utilities for allocating each job, and a linear constraint is applied to each job and each computing resource.
+Thus, DeDe can decompose the problem into $n$ per-job allocations and $m$ per-demand allocations.
 5.2 Traffic Engineering
-The traffic engineering problems [24, 54, 43, 42, 63] allocate traffic demand between datacenter nodes onto network links in a WAN topology.
-Maximize total flow.  The problem focuses on maximizing the total flow allocation.
+Traffic engineering problems [24, 55, 44, 43, 65] allocate traffic demands between datacenters onto the network links in a wide-area network (WAN) topology.
+Maximizing total flow.  This problem variant focuses on maximizing the total flow allocation.
   • Variables.
-The flow allocation from datacenter node pair $(s,t)$ onto network link $(u,v)$ is denoted as $x_{(u,v)(s,t)}\geq 0$.
-In common path-based traffic engineering, flows between node pair $(s,t)$ are allocated over links along pre-configured paths $P_{(s,t)}$ only.
+Let $x_{(u,v)(s,t)}\geq 0$ denote the amount of flow from a datacenter pair $(s,t)$ assigned to a network link $(u,v)$.
+In common path-based traffic engineering, flows between each node pair $(s,t)$ are allocated only over links along pre-configured paths $P_{(s,t)}$.
   • Objective.
-For a node pair $(s,t)$, the inflow it incurs to node $v$ is computed as $\text{inflow}(v,(s,t))=\sum_{(u,v)\in P_{s,t}}x_{(u,v)(s,t)}$ and outflow as $\text{outflow}(v,(s,t))=\sum_{(v,u)\in P_{s,t}}x_{(v,u)(s,t)}$.
-The total flow allocated for the node pair $(s,t)$ can be represented as $\text{inflow}(t,(s,t))$, which denotes the total inflow toward the destination $t$.
-Therefore, the objective of maximizing total flows between all node pairs can be expressed as the sum:
+For each node pair $(s,t)$, the inflow it incurs to node $v$ is computed as: $\text{inflow}(v,(s,t))=\sum_{(u,v)\in P_{s,t}}x_{(u,v)(s,t)}$, and outflow as: $\text{outflow}(v,(s,t))=\sum_{(v,u)\in P_{s,t}}x_{(v,u)(s,t)}$.
+Consequently, the total flow allocated for the node pair $(s,t)$ is captured by $\text{inflow}(t,(s,t))$, the total inflow toward the destination $t$.
+Therefore, the objective of maximizing the total flow between all node pairs can be expressed as:
     $\operatorname*{Maximize\leavevmode\nobreak\ }_{x}\sum_{(s,t)}\text{inflow}(t,(s,t)).$
-  • Resource constraints.  For each link $(u,v)$, the total flow allocation from all node pairs must not exceed its link capacity $c_{u,v}$:
+  • Resource constraints.  For each link $(u,v)$, the cumulative flow from all node pairs traversing it must not exceed its link capacity $c_{u,v}$:
     $\sum_{(s,t)}x_{(u,v)(s,t)}\leq c_{u,v},\quad\forall(u,v)\in E.$
@@
     $\begin{array}[]{l}0\leq\text{inflow}(t,(s,t))\leq d_{s,t},\\ \text{inflow}(v,(s,t))=\text{outflow}(v,(s,t)),\forall v\neq s,t,\\ \end{array}\quad\forall(s,t).$
-Min-max link utilization.
-The problem of min-max link utilization aims to allocate demands fairly over network links.
-The problem shares the same variables and linear demand constraints as the max total flow problem.
-  • Objective.  The utilization of a link $(u,v)$ is the ratio of the total flow assigned to the link to its capacity. Our objective is to minimize the max link utilization across the network:
+Minimizing max link utilization.
+This variant aims to evenly distribute traffic demands across network links to avoid overloading any single link.
+The utilization of link is commonly defined as the (uncapped) ratio of total flow assigned to a link to the link’s capacity, although it cannot actually exceed 100% in practice.
+This formulation shares the same variables and constraints as above, with a revised objective:
+  • Objective.  The goal is to minimize the max link utilization across the network:
     $\operatorname*{Minimize\leavevmode\nobreak\ }_{x}\max_{(u,v)\in E}\frac{\sum_{(s,t)}x_{(u,v)(s,t)}}{c_{u,v}}.$
-Using DeDe.
-The traffic engineering problem aligns the separable structure for DeDe: the objectives are the sum of allocation quality per demand or the maximum of allocation quality per link, respectively.
-Linear constraints are applied to each demand and each link.
+Applying DeDe.
+The traffic engineering problem aligns with the separable structure of DeDe: the objectives are expressed as either a sum over per-demand utilities or as a maximum over per-link utilities, and linear constraints are applied to each demand and each link.
 Therefore, DeDe can decompose the problem into $|E|$ per-link allocations and $|V|^{2}$ per-demand allocations for parallelism.
-5.2 Cluster Scheduling
-Cluster scheduling problems [61, 44, 43, 42] involves allocating jobs to computing resources (e.g., CPUs, GPUs) in heterogeneous clusters.
-Max-min allocation.  This problem aims to distribute the heterogeneous cluster resources fairly among jobs based on their effective throughput.
+In practice, due to the large overhead of maintaining $|V|^{2}$ per-demand subproblems, we group these subproblems by their sources, reducing the total number of subproblems to just $|V|$.
+5.3 Load Balancing
+Load balancing problems [14, 49, 53, 44] allocate data shards among storage servers to scale out query loads in a distributed store.
+Load balancing is a non-convex problem with integer variables, wherein DeDe can provide a high-quality solution empirically, as discussed in §4.2.
+Minimizing shard movements.  This problem aims to minimize shard movements across servers during load changes while keeping the load nearly balanced on each server.
   • Variables.
-With jobs time-sliced onto the available heterogeneous computing resources [44, 43], the allocation is a matrix $x\in[0,1]^{n\times m}$ where $x_{ij}$ is the fraction of time each job $j$ spent on each computing resource type $i$.
+The allocation plan is encoded in a matrix $x\in[0,1]^{n\times m}$, where $x_{ij}$ is the fraction of data shard $j$ assigned to server $i$.
+Additionally, a binary matrix $x^{\prime}$ captures shard placement, where $x^{\prime}_{ij}=1$ if the data shard $j$ is located on server $i$ (i.e., when $x_{ij}>0$); otherwise, $x^{\prime}_{ij}=0$.
   • Objective.
-Given an allocation matrix $x$, we adopt a normalized effective throughput of job $j$ from POP [41], which is a linear function on $x_{*j}$.
-The objective is to maximize the minimum allocation of normalized effective throughput:
-    $\operatorname*{Maximize\leavevmode\nobreak\ }_{x}\min_{j}\leavevmode\nobreak\ \text{throughput}(j,x_{*j}).$
-  • Resource constraints.  For each computing resource type $i$, the total time allocation cannot exceed the number of that computing resource type:
-    $\sum_{j}x_{ij}z_{j}\leq\text{num\_workers}_{i},\quad\forall i,$
-    where $z_{j}$ is the computing resources requested by job $j$.
-  • Demand constraints.  For each job $j$, The sum of all fractions of time must be within 1:
-    $\sum_{i}x_{ij}\leq 1,\quad\forall j.$
-Proportional fairness.
-The proportional fairness problem maximizes overall resource utilization while ensuring minimum service for each job.
-The problem shares the same variables, resource constraints, and capacity constraints.
-  • Objective.  Proportional fairness aims to maximize the sum of log utilities of each job:
-    $\operatorname*{Maximize\leavevmode\nobreak\ }_{x}\sum_{j}\log(\text{throughput}(j,x_{*j})).$
-Using DeDe.
-The cluster scheduling exhibits the separable structure required by DeDe: The objective function is a minimum of allocation quality per job and a sum of allocation quality per job, respectively.
-Additionally, linear constraints are applied to each job and each computing resource.
-Thus, DeDe can decompose the problem into $n$ per-job allocations and $m$ per-demand allocations.
-5.3 Load Balancing
-Load balancing problems [16, 48, 52, 43] allocate data shards among servers to scale out query loads in a distributed store.
-Minimize shard movements.  This problem aims to minimize shard movements across servers during load changes while keeping the load nearly balanced on each server.
-  • Variables.
-The allocation is a matrix $x\in[0,1]^{n\times m}$, where $x_{ij}$ is the fraction of data shard $j$ allocated to server $i$.
-Additionally, a binary matrix $x^{\prime}$ captures shard placement, where $x^{\prime}_{ij}=1$ if the data shard $j$ is on the server $i$, i.e., $x_{ij}>0$; otherwise $x_{ij}=0$.
-  • Objective.
-The movement of data shard $j$ allocated to server $i$ can be represent as $(1-T_{ij})x^{\prime}_{ij}f_{i}$, where $T_{ij}$ denotes the initial shard placement and $f_{j}$ denotes the memory footprint of data shard $j$.
-The objective of minimizing total shard movement can be expressed as the sum:
+The movement of data shard $j$ allocated to server $i$ can be represent as $(1-T_{ij})x^{\prime}_{ij}f_{i}$, where $T_{ij}$ denotes the initial shard placement, and $f_{j}$ denotes the memory footprint of data shard $j$.
+The objective of minimizing total shard movement can be expressed as:
     $\operatorname*{Minimize\leavevmode\nobreak\ }_{x}\sum_{j}\sum_{i}(1-T_{ij})x^{\prime}_{ij}f_{i}.$
-  • Resource constraints.  For each server $i$, the total query load must be close to the average query load $L$; The total memory usage must be within memory capacity:
+  • Resource constraints.  For each server $i$, the total query load must be close to the average query load $L$. Meanwhile, the total memory usage must be within the memory capacity:
     $\begin{array}[]{l}L-\epsilon\leq\sum_{j}x_{ij}l_{j}\leq L+\epsilon,\\ \sum_{j}x^{\prime}_{ij}f_{j}\leq\text{memory}_{i},\end{array}\quad\forall i,$
     where $l_{j}$ is the query load on data shard $j$.
-  • Demand constraints.  Each data shard $j$ must allocate all the query load among servers:
+  • Demand constraints.  Each data shard $j$ must be entirely allocated across servers:
     $\sum_{i}x_{ij}=1,\quad\forall j.$
-Using DeDe.
-Given the separable structure, where the objective is the sum of allocation quality per data shard and linear constraints apply to each server, and each shard, DeDe can decompose the problem into $n$ per-server allocations and $m$ per-shard allocations.
+Applying DeDe.
+The load balancing problem also demonstrates a separable structure: the objective is the sum of allocation quality per data shard, while linear constraints are applied independently to each server and each shard.
+This allows DeDe to decompose the problem into $n$ per-server allocations and $m$ per-shard allocations.
 6 Implementation of DeDe
 We have implemented DeDe as a Python package and uploaded it to PyPI for easy installation: pip install dede.
-Listing 1 showcases a simple resource allocation problem solved using our DeDe package.
-In line 4, we initially create a $N\times M$ non-negative allocation variable x, where x[i,j] represents the allocation of resource i to demand j.
-We then create a constraint for each resource in line 7–8 and a constraint for each demand in line 9–10, as well as an objective of minimizing the sum of x in line 13.
-Finally, we construct the problem in line 16–17 and solve it in line 18.
-1import dede
-2
-3# Create allocation variables
-4x = dede.Variable((N, M), nonneg=True)
-5
-6# Create the constraints
-7resource_constrs = [
-8 x[i,:].sum() <= 1 for i in range(N)]
-9demand_constrs = [
+Listing 1 showcases a basic resource allocation example using DeDe.
+Line 5 creates an $N\times M$ non-negative allocation matrix x, where x[i,j] represents the fraction of time that demand j runs on resource i.
+Lines 12–13 define a constraint for each resource using parameters, and Lines 14–15 define a constraint for each demand.
+The objective, defined in Line 18, is to maximize the sum of entries in x.
+Finally, the problem is constructed in Lines 21–22 and solved in Line 23.
+Built on top of the popular optimization modeling language cvxpy [15, 3], DeDe inherits most of its syntax and APIs, such as Variable(), Parameter(), and Maximize().
+A notable distinction is that DeDe requires users to explicitly separate resource constraints and demands constraints when initializing a problem (Line 22).
+As a parallel optimization tool, DeDe allows users to configure the number of CPUs (Line 23); by default, all available cores are used.
+Internally, DeDe solves the optimization in three stages:
+  • Problem parsing.
+DeDe transforms all inequality constraints into equivalent equality constraints through non-negative slack variables.
+E.g., x[:,j].sum() <= 1 is converted to x[:,j].sum() + sj == 1 with a slack variable sj.
+Then slack variables are treated just like other variables.
+  • Problem building.
+DeDe organizes resource constraints into disjoint per-resource groups and demand constraints into disjoint per-demand groups.
+For each group, a per-resource or per-demand subproblem is constructed with cvxpy following Equations 8 and 9.
+  • Problem solving.
+DeDe solves these subproblems in parallel using multiple CPU cores.
+During ADMM iterations, only the parameters are updated, avoiding the overhead of rebuilding problems in cvxpy.
+Similarly, for the same problem with varying resources and demands, only the relevant parameters are updated.
+1import numpy as np
+2import dede as dd
+3
+4# Create allocation variables
+5x = dd.Variable((N, M), nonneg=True)
+6
+7# Create parameters
+8param = dd.Parameter(
+9 N, value=np.random.uniform(0, 1, N))
+10
+11# Create constraints
+12resource_constrs = [
+13 x[i,:].sum() <= param[i] for i in range(N)]
+14demand_constrs = [
 15 x[:,j].sum() <= 1 for j in range(M)]
-11
-12# Create an objective
-13obj = dede.Minimize(x.sum())
-14
-15# Construct and solve the problem
-16prob = dede.Problem(
+16
+17# Create an objective
+18obj = dd.Maximize(x.sum())
+19
+20# Construct and solve the problem
+21prob = dd.Problem(
 22 obj, resource_constrs, demand_constrs)
-18prob.solve(num_cpus=64)
+23prob.solve(num_cpus=64, solver=dd.ECOS)
 Listing 1: Resource allocation example with DeDe package.
-DeDe borrows the interface from cvxpy [17, 4].
-DeDe inherits most of cvxpy’s methods, e.g.
-Variable(.) in line 4, Minimize(.) in line 13.
-One difference from cvxpy is that users must specify resource constraints and demands constraints separately (line 16–17) when initializing a problem.
-As a parallel solution, DeDe allows users to specify the number of CPU cores (line 18); if not specified, DeDe uses all available CPU cores.
-Behind DeDe’s interface, the resource allocation problem is solved in the following three steps.
-  • Problem parsing.
-Inequality constraints are converted into equality constraints with a non-negative slack variable s., e.g., inequality x[0,:].sum() <= 1 can be converted to equality x[0,:].sum() + s == 1.
-The slack variable can be regarded as an additional part of the problem variables and be optimized together.
-  • Problem building.
-Resource constraints are organized into independent per-resource subgroups without shared variables, while demand constraints are similarly organized into separate independent per-demand groups.
-For each subgroup, a per-resource or per-demand problem is built using cvxpy, with the new objective formulated based on the augmented Lagrangian method outlined in § 3.2.
-  • Problem solving.
-Per-resource or per-demand problems in cvxpy are solved in parallel across multiple cores.
-Between ADMM iterations, only the parameters are updated, avoiding the need to rebuild the entire cvxpy problem.
-Similarly, for the same problem with varying resources and demands, only the corresponding parameters are updated.
-We use Ray [40] to parallelize our computation across multiple cores.
-Python has limited parallelism in its multithreading implementation due to the notorious global interpreter lock (GIL).
-Instead, we choose Ray, which provides a wrapper for multiple Python interpreter processes, so as to avoid GIL.
-We also benefit from Ray’s production-grade implementation of inter-process coordination mechanisms to avoid building them from scratch using Python’s built-in low-level multiprocessing library.
+We use Ray [41] to parallelize our computation across multiple cores, as Python’s native multithreading model is constrained by the (notorious) global interpreter lock (GIL).
+Ray circumvents this issue by managing multiple Python interpreter processes, effectively bypassing the GIL.
+Moreover, Ray offers robust support for inter-process communication, sparing us from manually implementing such mechanisms using Python’s lower-level multiprocessing primitives.
 7 Evaluation
@@
   2. 2.
-    How do DeDe and state-of-the-art approaches react to changes in problem granularity, temporal fluctuation, spatial redistribution, and link failures? (§7.2)
+    How do DeDe and baseline approaches react to changes in problem granularity, temporal fluctuation, spatial redistribution, and link failures? (§7.2)
   3. 3.
-    What is the contribution of each design in DeDe? (§7.3)
-We focus on four approaches throughout our evaluation:
+    What is the individual contribution of each design component in DeDe? (§7.3)
+Our evaluation compares four approaches:
   • Exact sol.
-Exact Sol solves the original resource allocation problem using commercial optimization solvers.
-We reuse the implementation from POP, which employs Gurobi [22], cvxpy [17, 4], and CPLEX [15] for traffic engineering, cluster engineering, and load balancing, respectively.
+This baseline solves the original resource allocation problem using commercial solvers.
+We adopt the implementation from POP, which utilizes cvxpy [15, 3], Gurobi [22], and CPLEX [25] for cluster engineering, traffic engineering, and load balancing, respectively.
   • POP.
-POP-$k$ [43] randomly splits the problem into $k$ smaller subproblems, applies commercial solvers in each, and coalesces the resulting $k$ sub-allocations into a global allocation.
-However, POP only simulates the parallel execution by solving subproblems sequentially, collecting the solving time of each, and calculating the parallel solving time mathematically.
-  • DeDe.  DeDe is our truly parallel implementation, with details described in Section 6.
+POP-$k$ [44] randomly splits the problem into $k$ smaller subproblems, applies commercial solvers to each, and coalesces the resulting $k$ sub-allocations into a global allocation.
+However, POP only simulates the parallel execution by solving subproblems sequentially and calculating the parallel solving time mathematically.
+  • DeDe.  DeDe is our truly parallel implementation (§6). By default, the solution from the previous optimization interval is used to warm-start the subsequent interval.
   • DeDe.
-To facilitate a fair comparison with POP, we create a variant of DeDe, denoted as DeDe, following POP-style simulation where subproblems are solved sequentially and a parallel solving time is calculated mathematically.
+For a fair comparison with POP, we create a variant of DeDe, denoted as DeDe, following POP’s simulation methodology.
+DeDe solves subproblems sequentially and estimates the parallel solving time mathematically.
 We also consider domain-specific approaches apart from these four approaches.
-In traffic engineering, we evaluate a demand-pinning approach (“Pinning” for short) similar to a prior work [41], with the top 10% of demands allocated using optimization engines and the remaining assigned to the shortest paths.
-We also evaluate Teal [63], a learning-accelerated traffic engineering approach with massive parallelism achieved on the GPU.
 In cluster scheduling and load balancing, we evaluate the greedy heuristic algorithms, Gandiva [63] and E-Store [53], respectively.
-We compare these approaches based on their allocation quality and computation times.
-The computation time is measured on 64 CPU cores (2$\times$ Intel Xeon Gold 6142); an additional GPU (Nvidia Titan RTX) is made available for Teal.
+In traffic engineering, we evaluate a demand-pinning approach (“Pinning”) similar to a prior work [42], where the top 10% of demands are allocated using optimization engines and the rest are assigned to shortest paths.
+We also evaluate Teal [65], a learning-accelerated traffic engineering approach that achieves massive parallelism using a GPU.
+All approaches are evaluated in terms of allocation quality and computation time.
+The computation time is measured using 64 CPU cores (2$\times$ Intel Xeon Gold 6142), with an additional GPU (Nvidia Titan RTX) available for Teal.
 7.1 DeDe vs. the State of the Art
-In this section, we compare DeDe against state-of-the-art approaches on resource allocation problems from Section 5.
+In this section, we compare DeDe against state-of-the-art approaches on resource allocation problems from §5.
+7.1.1 Cluster Scheduling
+We stress-test DeDe’s scalability in a large, heterogeneous cluster scheduling environment, where jobs are allocated in a time-sliced manner across 456 different types of compute resources, e.g., GPU/CPU instances that vary in vendor, generation, memory capacity, and other specifications.
+The quantity of each resource type is randomly drawn from $\{8,16,24,\ldots,64\}$, resulting in a total of 16,520 instances.
+We utilize the simulator in Gavel [45] and model job arrivals as a Poisson process with an average inter-arrival of 100 seconds.
+We update job allocation decisions every 6 minutes, repeating this process for 200 scheduling rounds (20 hours).
+For each job, the number of requested resource instances per type is drawn from $\{1,2,4,8,16,32\}$, and following prior work [59], 33% of the jobs are restricted to specific resource types.
+Job throughput is derived from relevant benchmarks [26, 36, 19].
+Please refer to Appendix A for more details on this setup.
+Figure 4: Results of the cluster scheduling variant that maximizes the minimum job throughput (max-min allocation).
+Max-min allocation.
+Figure 4 compares DeDe against other approaches in terms of max-min allocation.
+DeDe quickly attains a near-optimal max-min allocation of 0.94 in just 3 seconds and 0.99 within 10 seconds, demonstrating a better balance between max-min allocation and computation time compared with Exact sol., Gandiva, and POP.
+In contrast, Exact sol.
+experiences a significantly slower computation time of 156 seconds due to its sequential nature.
+The greedy heuristic, Gandiva, performs poorly, achieving only a max-min allocation of 0.43 (in 1.4 seconds).
+POP-4 provides a comparable tradeoff between time and accuracy to DeDe, but is 1.6$\times$ slower.
+The faster variant, POP-16, splits the problem into 16 smaller subproblems and achieves a lower max-min allocation of 0.9 (in 3.1 seconds), since smaller subproblems limit the ability to select optimal GPU/CPU assignments.
+DeDe, representing the same algorithm under idealized conditions similar to POP’s simulation, achieves a normalized max-min allocation of 0.99 and is 3.3$\times$ faster than DeDe in Figure 4.
+The reason is that DeDe executes each iteration more quickly and completes more iterations within a given time.
+The discrepancy in iteration time between DeDe and DeDe arises from three factors: (1) DeDe measures the end-to-end time, including problem solving, compilation, solution unpacking, and other associated tasks, while DeDe only accounts for the core solving time; (2) DeDe’s parallel implementation incurs overhead from cache contention, which slows computation; (3) In DeDe, each subproblem is statically pre-assigned to one of the processes, making it susceptible to straggler delays.
+In contrast, DeDe assumes perfect dynamic scheduling.
+Figure 5: Results of the cluster scheduling variant that maximizes proportional fairness among job throughputs.
+Proportional fairness.
+Figure 5 compares DeDe against other approaches in maximizing proportional fairness.
+This problem is particularly challenging due to the logarithmic form of the objective function.
+Exact sol., which uses the SCS solver in cvxpy, fails to reach optimality even after 5 hours of computation.
+As a result, several methods report normalized fairness scores that exceed 1 when normalized relative to Exact sol.
+Both DeDe and DeDe achieve greater speedups than POP by effectively breaking the problem into smaller and easier subproblems.
+Notably, DeDe and DeDe both reach a normalized proportional fairness of over 1 within 100 seconds.
+In contrast, POP-4 and POP-16 require 3,053 seconds and 682 seconds, respectively, to achieve comparable allocation quality, while POP-64 yields a significantly lower fairness score of only 0.31.
 7.1.2 Traffic Engineering
-Figure 3: Comparison for the problem of maximizing total flow in traffic engineering.
-We adopt the largest network in Teal’s dataset for traffic engineering problems.
-This 1,739-node network is adapted from an AS-level internet topology [12] for WAN purposes, with the traffic data collected from a global cloud provider and then mapped into the large network.
+Figure 6: Results of the traffic engineering problem variant that maximizes the total flow.
+We adopt the evaluation setup in Teal [65], with a focus on its largest topology—a 1,739-node network adapted from an internet topology for WAN purposes.
+The traffic matrices are collected from the production WAN of a global cloud provider and mapped onto the above topology.
 Maximize total flow.
-Figure 3 compares DeDe to state-of-the-art methods in maximizing total flow.
-DeDe achieves 90.8% satisfied demand within 30 seconds and 92% within 60 seconds, demonstrating a superior trade-off between satisfied demand and computation time compared to Exact Sol, POP, and Pinning.
-While POP-4 also reaches 92% satisfied demand, its average computation time is significantly longer at 1658 seconds.
+Figure 6 compares DeDe to the baselines in maximizing total flow.
+DeDe achieves 90.8% satisfied demand within 30 seconds and 92% within 60 seconds, showing a superior trade-off between satisfied demand and computation time compared with Exact sol., POP, and Pinning.
+Its idealized variant, DeDe, achieves a satisfied demand that is 8% higher than DeDe within 10 seconds.
+While POP-4 also reaches 92% satisfied demand, its average computation time is significantly longer, at 1,658 seconds.
 POP-16 and POP-64 improve parallelism by dividing the workload into more subproblems, reducing computation times to 456 seconds and 380 seconds, respectively.
-Note that the speedup of POP-64 is limited compared to POP-16 as the number of cores per subproblem drops from 4 to 1.
+Note that the speedup of POP-64 is limited compared with POP-16 as the number of cores per subproblem drops from 4 to 1.
 This increased parallelism also comes at a cost, with satisfied demand decreasing to 87.4% for POP-16 and 81.6% for POP-64.
@@
 As a result, it achieves a modest speedup, requiring 149 seconds to reach 87.3% satisfied demand.
-Teal, a learning-accelerated approach, leverages massive GPU parallelism to achieve 89% satisfied demand in just 1 second.
+Teal, a learning-accelerated approach, leverages massive GPU parallelism to achieve 89% satisfied demand in only 1 second.
 However, it relies on a highly customized machine-learning framework, demanding significant human effort for design and training.
-DeDe, representing the same algorithm under ideal conditions as POP, achieves a satisfied demand that is 8% higher than DeDe within 10 seconds in Figure 3 and consistently delivers superior allocation quality across the other figures.
-The reason is that DeDe executes each iteration more quickly, allowing it to complete more iterations within a given time, resulting in higher allocation quality.
-The difference in iteration time between DeDe and DeDe arises from three factors, as discussed in more detail in § 7.3: (1) DeDe measures the end-to-end time, including problem-solving, compilation, solution unpacking, and other associated tasks, while DeDe only takes the problem-solving time into account.
-(2) DeDe’s parallel implementation (§ 6) incurs overhead from cache contention, which slows down the problem-solving process.
-(3) In DeDe, each subproblem is statically pre-assigned to one of the processes, making it susceptible to stragglers.
-In contrast, DeDe assumes perfect dynamic scheduling.
-Figure 4: Comparison for the problem of minimizing max link utilization in traffic engineering.
+Figure 7: Results of the traffic engineering problem variant that minimizes the max link utilization.
+Figure 8: Results of the load balancing problem where the objective is to minimize shard movements.
 Min-max link utilization.
 Figure 7 compares DeDe with other approaches in minimizing maximum link utilization.
-DeDe achieves a low maximum link utilization of 1.67 within 10 seconds and 1.71 within 5 seconds.
-Exact Sol reaches 1.63 with an average computation time of 35 seconds, and it is faster than itself in maximizing total flow because minimizing max link utilization is "easier" and requires fewer iterations.
-Pinning takes 42.5 seconds, making it slower than DeDe and slightly slower than Exact Sol, as it requires additional time to rebuild the problem when the top 10% of node pairs change.
-POP’s variants, POP-4, POP-16, and POP-64, achieve a maximum link utilization of 1.70, 1.77, and 1.95 with 9, 2, and 0.4 seconds on average, respectively.
+Here, the link utilization metric serves as proxy for network congestion and is allowed to exceed 100% during optimization.
+DeDe attains a maximum link utilization of 1.67 within 10 seconds and 1.71 within 5 seconds.
+Exact sol.
+reaches 1.63 with a longer computation time of 35 seconds.
+This is faster than maximizing total flow because minimizing max link utilization is computationally less challenging and thus requires fewer iterations.
+Pinning takes 42.5 seconds, making it slower than DeDe and slightly slower than Exact sol., as it requires additional time to rebuild the problem when the top 10% of node pairs change.
+POP’s variants, POP-4, POP-16, and POP-64, achieve a maximum link utilization of 1.70, 1.77, and 1.95 within 9, 2, and 0.4 seconds on average, respectively.
 Although POP-16 is faster than DeDe, a fairer comparison between POP-16 and DeDe, both simulating parallelism, reveals that DeDe achieves a comparable computation time to POP-16.
-The domain-specific solution Teal takes only 0.3 seconds with massive parallelism from GPU with a maximum link utilization of 1.74.
-7.1.2 Cluster Scheduling
-We evaluate a large cluster scheduling problem of allocating jobs to 5000 GPU/CPUs of 196 different types.
-We collect real throughput from two computing benchmarks [25, 35].
-The number of GPU/CPUs for each type and the required number in each job is generated based on the distribution observed in traces from real GPU-sharing clusters [58].
-To reduce simulation time, we capture snapshots through the scheduling process rather than executing the entire simulation from start to finish.
-Figure 5: Comparison for the min-max allocation problem in cluster scheduling.
-Max-min allocation.
-Figure 5 compares DeDe against other approaches with the objective of max-min allocation.
-DeDe achieves a near-optimal max-min allocation of 0.97 within 5 seconds and 0.98 within 10 seconds, while DeDe reaches 0.98 within 2 seconds.
-Due to their sequential nature, Exact Sol and Gandiva encounter a slow computation time of 86 and 22 seconds.
-The greedy heuristic Gandiva even experiences a 0.56 low max-min allocation.
-POP-4 deliverers a comparable time-accuracy tradeoff as DeDe but is 3.4$\times$ slower than DeDe, which is simulated in the same way as POP did for a fair comparison.
-A faster POP variant, POP-16, exhibits a lower max-min allocation of 0.94, as it becomes more difficult for jobs to assess the ideal GPU/CPU among the limited selection in smaller subproblems.
-Figure 6: Comparison for the problem of maximizing proportional fairness in cluster scheduling.
+The domain-customized Teal takes only 0.3 seconds on a GPU to attain a maximum link utilization of 1.74.
+7.1.3 Load Balancing
 (a) Granularity changes.
@@
 Figure 9: Normalized satisfied demand in traffic engineering against granularity, temporal, and spatial changes.
-Figure 8: Comparison for the problem of minimizing shard movements in load balancing.
-Proportional fairness.
-Figure 6 compares DeDe against other approaches in maximizing proportional fairness.
-As the optimization problem becomes more difficult with the convex logarithm function in the objective, both DeDe and DeDe achieve a greater speedup than POP by thoroughly breaking down the problem into smaller and easier pieces.
-DeDe and DeDe reach a proportional fairness score of over 1,000 in 300 seconds and 100 seconds, respectively.
-In contrast, POP-4 takes 1,825 seconds to reach the same score, while POP-16 and POP-64 remain below 900 despite their shorter computation times.
-Exact Sol, though accurate, is prohibitively slow, requiring 12,432 seconds.
-7.1.3 Load Balancing
-We evaluate a load balancing problem of assigning 2048 data shards to 256 servers. In each round, we compute a new allocation of shards to servers given the new query loads [43].
+We follow the load balancing settings in POP [44], except for increasing the scale to 2,048 data shards distributed across 256 servers.
+In each round, a new shard-to-server allocation is computed based on updated query loads.
+We simulate for 100 rounds, using the first 20 as a warm-up.
+To ensure the problem’s feasibility, we set the tolerance parameter $\epsilon$ to 0.1.
 Minimize shard movements.
-Figure 8 compares DeDe with state-of-the-art approaches in minimizing shard movements.
-DeDe achieves an average of 24.1 shard movements in 15 seconds, slightly better than POP-4’s 25.4 shard movements within 16.9 seconds.
-In faster scenarios, DeDe achieves 35.1 shard movements in 2 seconds, while POP-4 achieves slightly fewer movements of 32.2 in 1.7 seconds.
-In a more fair comparison with POP, DeDe consistently costs 15–18% fewer shard movements than POP given a similar time limit.
-Exact Sol achieves the minimum shard movements of 13.2 but is significantly slower, taking an average of 1,572 seconds due to the NP-hardness of solving the mixed-integer linear programming problem inherent to load balancing.
-On the other hand, the greedy heuristic represents the opposite end of the time-accuracy trade-off, completing in just 25 milliseconds but requiring 50 shard movements.
+Figure 8 compares DeDe with baselines in minimizing shard movements.
+DeDe achieves an average of 20.1 shard movements in 15 seconds, better than POP-4’s 21.5 shard movements within 133 seconds.
+In faster scenarios, DeDe achieves 25.4 shard movements in 5 seconds, while POP-16 achieves an average movement of 26 in a slighter less time of 3.7 seconds.
+For a fair comparison, DeDe consistently achieves around 16% fewer shard movements than POP, given a similar time limit.
+Exact sol.
+requires 20.9 shard movements on average, but is significantly slower with an average of 4,820 seconds.
+This slower computation is due to the NP-hard nature of solving the mixed-integer linear program underlying load balancing.
+Notably, Exact sol.
+does not produce the minimum possible shard movements because optimizing each scheduling round independently does not guarantee an optimal solution across the entire series.
+At the opposite end of the time-accuracy trade-off, the greedy heuristic provides a rapid but less accurate solution, completing in just 2 milliseconds but averaging 73 shard movements after naively fixing its constraint violations.
 (a) Speedup when varying CPU cores.
 (b) Convergence rate.
-(c) Comparison to more optimization methods.
+(c) Comparing different optimization methods.
 Figure 10: Micro-benchmarks with the objective of maximizing total flow in traffic engineering.
@@
 7.2 Robustness
-We evaluate the robustness of DeDe and other baselines against changes in problem granularity, temporal fluctuations, spatial redistribution, and link failures.
-In all these evaluations, DeDe is run for 30 seconds.
-To mitigate the impact of the changing resources and demands, we normalize the satisfied demand by the optimal satisfied demand from Exact Sol.
+Using traffic engineering as an example, we evaluate the robustness of DeDe and other baselines against changes in problem granularity, temporal dynamics, spatial redistribution, and link failures.
+In these experiments, DeDe is run for 30 seconds, and we normalize the satisfied traffic demand by the optimal one achieved by Exact sol.
 Changes of problem granularity.
-The granular problem required by POP, where resources are mutually interchangeable, is impractical in reality.
-In traffic engineering, for example, demands are constrained to capacity resources on pre-configured paths.
+The granular problem structure required by POP, where each demand only requests a small fraction of interchangeable resources, can be impractical.
+In traffic engineering, for example, demands are constrained by the link capacity on specific pre-configured paths.
 Similarly, in cluster scheduling, 33% of GPU tasks in production clusters are limited to specific GPU types [59].
-Using traffic engineering as an example, we evaluate all approaches across varying levels of problem granularity, as depicted in Figure 7(a).
+In Figure 9(a), we present results on varying levels of problem granularity using traffic engineering as an example.
 To quantify resource interchangeability, we use the mean edge betweenness centrality [9, 10], which measures the average percentage of demands served by a given edge.
@@
 This is because the reduced edge selection simplifies allocation decisions.
-Temporal fluctuation.
-We introduce temporal fluctuations [63] to the traffic matrices for robustness assessment.
+Temporal dynamics.
+Next, we introduce temporal fluctuations [65] to the traffic matrices for robustness assessment.
 For each demand, we calculate the variance $\sigma^{2}$ in its changes between consecutive time slots and create a new normal distribution $N(0,k\sigma^{2})$.
@@
 Pinning’s performance decreases by 8.3%, as its strategy of focusing only on top demands relies on the presence of a heavy-tailed demand distribution.
-Teal has a slight decrease up to 2.4% due to unseen demand distribution in training.
-This is a relatively smaller decrease than the temporal fluctuation, as Teal can still handle simple, evenly distributed traffic demands.
+Teal has a slight decrease of up to 2.4% due to unseen demand distribution in training.
+This is a relatively smaller decrease than the temporal fluctuation, because Teal can still handle simple, evenly distributed traffic demands.
 POP has a slight increase of 1.4% as demands are more granular, where no demand requires a significant fraction of the total resources.
 Link failures.
-In practice, large-scale inter-datacenter link failures are exceedingly rare.
-Even in the event of fiber link failures, they typically do not result in capacity loss on an inter-datacenter link, unless caused by a fiber cut [51].
-As a stress test, however, we assess extreme failure scenarios described in previous work [67] and deliberately introduce 50, 100, and 200 link failures.
-Figure 10 illustrates the offline satisfied demand following these link failures, and we observe a consistent trend across different link failure scenarios across all approaches.
-Specifically, all approaches drop 0.6–1.1% under 50 link failures, 1.5–2.1% under 100 link failures, and 3–4.2% under 200 link failures.
-The reason is that these link failures constitute only a minor portion of the total 8,558 links, allowing all approaches to exhibit robustness given adequate time for re-computations.
+Widespread failures across inter-datacenter links are relatively rare in real-world operations unless triggered by physical fiber cuts [52].
+However, we simulate severe failure scenarios, introducing 50, 100, and 200 simultaneous link failures.
+As shown in Figure 11, the satisfied demand declines consistently across all methods.
+Specifically, all approaches drop by 0.6–1.1% under 50 link failures, 1.5–2.1% under 100 link failures, and 3–4.2% under 200 link failures.
+These drops stem from the fact that the failed links represent only a small portion of the total 8,558 links, allowing all approaches to recompute flow allocation given adequate time.
 7.3 Micro-benchmarks
-We conduct micro-benchmarks to evaluate the impact of DeDe’s design components, with all evaluations performed in maximizing total flow in traffic engineering.
+We conduct micro-benchmarks to evaluate the impact of DeDe’s design components, with all experiments performed in maximizing total flow in traffic engineering.
 Speedup when varying CPU cores.
-Figure 9(a) evaluates the speedup of DeDe, DeDe, and Exact Sol with 1, 4, 16, and 64 CPU cores.
+Figure 10(a) evaluates the speedup of DeDe, DeDe, and Exact sol.
+with 1, 4, 16, and 64 CPU cores.
 DeDe achieves 61.7$\times$ speedup under 64 cores, where the slight gap occurs as we cannot exactly divide the total computation time evenly across 64 cores.
@@
 However, this linearity diminishes with 64 cores, where DeDe only achieves 18.2$\times$ speedup.
-This is because utilizing all 64 cores for parallelism incurs overhead from cache contention and stragglers, as discussed in Section 7.1.
-In contrast to the linear speedup of DeDe and DeDe, the speedup of Exact Sol is sublinear and marginal, achieving only 3.4$\times$ speeds with 64 CPU cores.
+This is because utilizing all 64 cores for parallelism incurs overhead from cache contention and stragglers, as discussed in §7.1.1.
+In contrast to the linear speedup of DeDe and DeDe, the speedup of Exact sol.
+is sublinear and marginal, achieving only 3.4$\times$ speeds with 64 CPU cores.
 This is because of commercial solvers’ sequential nature, e.g., the simplex method [46] in linear programming requires thousands to millions of small steps to reach the optimal solution.
 Convergence rate.
-Both Teal and DeDe employ ADMM for parallel optimization, but their approaches differ.
-Teal’s ADMM adapts its computations based on matrix multiplication to leverage GPU calculation capabilities.
-Figure 9(b) illustrates the convergence rates of Teal and DeDe under various initialization.
-DeDe takes 30 seconds to surpass 90% satisfied demand.
-However, when initialized with Teal’s coarse neural network solution, DeDe achieves this within 20 seconds.
-Conversely, with naive initialization, DeDe requires more than 60 seconds.
-Teal, benefiting from the massive parallelism of GPUs, reaches 90% satisfied demand within 10 seconds.
-However, its adapted ADMM limits its performance, capping its satisfied demand at 91% at 60 seconds compared to DeDe’s 92%.
-Comparison to more optimization methods.
-Figure 9(c) investigates more constrained optimization methods other than DeDe’s ADMM method in maximizing the total flow in traffic engineering.
-Section 7.1 has covered the computation time of the optimal Exact sol, the suboptimal POP-64, and the near-optimal DeDe.
-As DeDe replaces constraints with additional terms in the objective function, we compare similar methods, including the penalty method and the augmented Lagrangian method.
-The augmented Lagrangian method is the same as in DeDe’s ADMM but without decoupling and decomposition, and thus it is over 3$\times$ slower than DeDe in reaching a satisfied demand of over 90%.
-The penalty method, which gradually increases penalty parameters from 0 to infinity to enforce the converted constraints in the objective function, converges slowly and is over 30$\times$ slower than DeDe in reaching a satisfied demand of over 90%.
+Teal generates a coarse, initial solution to traffic engineering using neural networks and fine-tunes it through ADMM.
+While Teal also “decouples” constraints, it does not further “decompose” the problem as DeDe does.
+Instead, it analytically and approximately minimizes a problem-specific augmented Lagrangian, directly projecting the result onto the non-negative feasible region.
+As a specialized method, Teal converges rapidly, achieving 90% satisfied demand within 10 seconds, as shown in Figure 10(b).
+In comparison, DeDe requires $\sim$30 seconds to attain comparable performance.
+In each optimization interval, DeDe is warm-started using the previous solution by default.
+When initialized instead with Teal’s neural network output, DeDe converges at a similar rate.
+In contrast, employing a naive initialization—where each demand is equally split across available paths—reduces DeDe’s convergence speed by nearly half.
+Alternative optimization methods.
+Figure 10(c) evaluates alternative constrained optimization techniques for maximizing total flow in traffic engineering—the penalty method [4] and the augmented Lagrangian method [23], comparing them with the ADMM-based approach used by DeDe.
+Both methods aim to solve the reformulated optimization problem in Equation 3.1 by jointly (and naively) optimizing variables $x$ and $z$, but they differ in how they enforce constraints.
+The penalty method introduces penalty terms into the objective function to discourage constraint violations.
+It starts with a small penalty coefficient and gradually increases it toward infinity, thereby driving the solution toward feasibility.
+However, this process typically involves solving a series of increasingly ill-conditioned problems, resulting in slow convergence.
+In our evaluation, it was more than 30$\times$ slower than DeDe in achieving a satisfied demand of over 90%.
+The augmented Lagrangian method improves upon the penalty method by combining both penalty terms and Lagrange multipliers.
+This helps enhance convergence stability and typically reduces the number of required iterations.
+However, unlike the ADMM approach used in DeDe, it lacks the benefit of decomposition and parallelism.
+As a result, despite its faster convergence over the basic penalty method, it remains over 3$\times$ slower than DeDe in reaching a satisfied demand of over 90%.
 8 Related Work
@@
 Shoofly [52] allocates light wavelengths of optical fiber to network shortcuts abstracted from optical bypasses, in order to minimize the hardware costs of provisioning long-haul capacity.
+By leveraging the commonly observed separable structure of real-world resource allocation problems, DeDe is applicable to the aforementioned cases as well as a broad range of others.
 Optimization algorithms.
@@
 In mixed-integer linear programming like load balancing, the branch-and-bound method [7] divides the problem into smaller subproblems, or branches, and then eliminates certain branches based on bounds on the optimal solution at each iteration.
-For more complex optimization problems, such as cone programming when maximizing proportional fairness in cluster scheduling, the ECOS algorithm [18] adopts a similar iterative process as the barrier methods, with search directions found by a symmetric indefinite KKT system.
+For more complex optimization problems, such as cone programming when maximizing proportional fairness in cluster scheduling, the ECOS algorithm [16] adopts a similar iterative process as the barrier methods, with search directions found by a symmetric indefinite KKT (Karush-Kuhn-Tuck) system.
 All these constrained optimization algorithms operate sequentially and iteratively, and thus introduce bottlenecks in solving time.
+In contrast, DeDe decouples and decomposes the problem into many parallel subproblems, and thus allows faster problem solving.
 Decomposition for scalability.
@@
 [17] employ the Benders decomposition algorithm to parallelize the design of a specialized cross-layer backbone network, offering performance guarantees for a particular allocation problem.
+While these prior decompositions rely on restrictive assumptions or problem-specific designs, DeDe offers a general and theoretically grounded approach solution for diverse resource allocation problems.
 9 Conclusion
-In this paper, we scale real-world resource allocation problems from a fresh perspective, observing that most of them exhibit an inherent separable structure.
+In this paper, we scale real-world resource allocation problems from a different lens, observing that the vast majority of them exhibit an inherent separable structure.
 This key insight allows us to systematically decouple resource and demand constraints and decompose the original optimization problem into independent and parallel subproblems for allocating individual resources and demands.
-We implement this approach, DeDe, as a library to augment a widely-used, open-source solver (cvxpy), offering a push-button solution that accelerates large-scale resource allocation.
+We implement this approach, DeDe, as a Python package, offering a push-button solution that accelerates large-scale resource allocation.
 Through extensive evaluation across diverse resource allocation tasks, we demonstrate that DeDe significantly improves scalability while closely approximating the quality of exact solutions.
-Appendix A Mathematical Representation of DeDe
-A resource allocation problem can be written as
-$\min_{x\in\mathcal{X}}\quad$ $\sum_{i}f_{i}(x_{i*})+\sum_{j}g_{j}(x_{*j}),$ (1)
-$s.t.\quad$ $R_{i}x_{i*}=r_{i},\quad\forall i,$ (2)
-$D_{j}x_{*j}=d_{j},\quad\forall j,$ (3)
-where Equation 1 is the objective, and Equation 2 and Equation 3 are resource constraints and demand constraints, respectively.
-Our first step decouples resource and demand constraints.
-We create a copy of the allocation variable $x$, denoted as $z$, and replace the variable $x$ in demand constraints (Equation 3) as well as the demand-related portion of the objective function (Equation 1) with the variable $z$.
-We also add a new constraint $x=z$ (Equation 6) to keep $x$ and $z$ equal.
-$\min_{x\in\mathcal{X}}\quad$ $\sum_{i}f_{i}(x_{i*})+\sum_{j}g_{j}(z_{*j}),$ (1)
-$s.t.\quad$ $R_{i}x_{i*}=r_{i},\quad\forall i,$ (2)
-$D_{j}z_{*j}=d_{j},\quad\forall j,$ (3)
-$x-z=0.$ (6)
-We rewrite our problem in the Augmented Lagrangian to move the constraints to the objective function:
-$\begin{split}&\mathcal{L}_{\rho}(x,z,\alpha,\beta,\lambda)=\sum_{i}f_{i}(x_{i*})+\sum_{j}g_{j}(z_{*j})\\ &+\frac{\rho}{2}\Bigl{(}\sum_{i}||R_{i}x_{i*}-r_{i}+\alpha_{i}||_{2}^{2}+\sum_{j}||D_{j}z_{*j}-d_{j}+\beta_{j}||_{2}^{2}\\ &+||x-z+\lambda||_{F}^{2}\Bigr{)}-\frac{\rho}{2}(||\alpha||_{2}^{2}+||\beta||_{2}^{2}+||\lambda||_{F}^{2}).\end{split}$
-ADMM allows partial optimization alternatively on $x$ and $z$:
-$x^{(k+1)}$ $:=\operatorname*{arg\,min}_{x\in\mathcal{X}}\mathcal{L}_{\rho}(x,z^{(k)},\alpha^{(k)},\beta^{(k)},\lambda^{(k)})$ (7)
-$z^{(k+1)}$ $:=\operatorname*{arg\,min}_{z}\mathcal{L}_{\rho}(x^{(k+1)},z,\alpha^{(k)},\beta^{(k)},\lambda^{(k)})$ (8)
-$\alpha_{i}^{(k+1)}$ $:=\alpha_{i}^{(k)}+R_{i}x^{(k+1)}_{i*}-r_{i},\forall i$
-$\beta_{j}^{(k+1)}$ $:=\beta_{j}^{(k)}+D_{j}z^{(k+1)}_{*j}-d_{j},\forall j$
-$\lambda^{(k+1)}$ $:=\lambda^{(k)}+x^{(k+1)}-z^{(k+1)}$
-Our second step further decomposes the partial optimization into optimizing each row in $x$ and each column in $z$. We rewrite Equation 7 as
-$x^{(k+1)}=$ $\operatorname*{arg\,min}_{x\in\mathcal{X}}\mathcal{L}_{\rho}(x,z^{(k)},\alpha^{(k)},\beta^{(k)},\lambda^{(k)})$ (7)
-$=$ $\operatorname*{arg\,min}_{x\in\mathcal{X}}\sum_{i}f_{i}(x_{i*})+\frac{\rho}{2}\sum_{i}\big{\|}R_{i}x_{i*}-r_{i}+\alpha_{i}^{(k)}\big{\|}_{2}^{2}$
-$+\frac{\rho}{2}\big{\|}x-z^{(k)}+\lambda^{(k)}\big{\|}_{F}^{2}.$
-We rearrange the parentheses in Equation 2 to make the variables in each parenthesis mutually exclusive:
-$x^{(k+1)}=$ $\operatorname*{arg\,min}_{x\in\mathcal{X}}\sum_{i}\Bigl{(}f_{i}(x_{i*})+\frac{\rho}{2}\big{\|}R_{i}x_{i*}-r_{i}+\alpha_{i}^{(k)}\big{\|}_{2}^{2}$
-$+\frac{\rho}{2}\big{\|}x_{i*}-z_{i*}^{(k)}+\lambda_{i*}^{(k)}\big{\|}_{2}^{2}\Bigr{)}.$
-With the variable in each parenthesis mutually exclusive, we decompose to optimize per-resource allocation $x_{i*}$ individually:
-$\begin{split}x_{i*}^{(k+1)}=&\operatorname*{arg\,min}_{x_{i*}\in\mathcal{X}_{i*}}\Bigl{(}f_{i}(x_{i*})+\frac{\rho}{2}\big{\|}R_{i}x_{i*}-r_{i}+\alpha_{i}^{(k)}\big{\|}_{2}^{2}\\ &+\frac{\rho}{2}\big{\|}x_{i*}-z_{i*}^{(k)}+\lambda_{i*}^{(k)}\big{\|}_{2}^{2}\Bigr{)}.\end{split}$
-We follow a similar decomposing procedure for demand-constrained optimization on $z$:
-$z^{(k+1)}=$ $\operatorname*{arg\,min}_{z}\mathcal{L}_{\rho}(x^{(k+1)},z,\alpha^{(k)},\beta^{(k)},\lambda^{(k)})$ (8)
-$=$ $\operatorname*{arg\,min}_{z}\sum_{j}g_{j}(z_{*j})+\frac{\rho}{2}\sum_{j}\big{\|}D_{j}z_{*j}-d_{j}+\beta_{j}^{(k)}\big{\|}_{2}^{2}$
-$+\frac{\rho}{2}\big{\|}x^{(k+1)}-z+\lambda^{(k)}\big{\|}_{F}^{2}$
-$=$ $\operatorname*{arg\,min}_{z}\sum_{j}\Bigl{(}g_{j}(z_{*j})+\frac{\rho}{2}\big{\|}D_{j}z_{*j}-d_{j}+\beta_{j}^{(k)}\big{\|}_{2}^{2}$
-$+\frac{\rho}{2}\big{\|}x^{(k+1)}_{*j}-z_{*j}+\lambda_{*j}^{(k)}\big{\|}_{2}^{2}\Bigr{)}.$
-The optimization of $z$ can now be decomposed into optimization on per-demand allocation $z_{*j}$:
-$\begin{split}z_{*j}^{(k+1)}=&\operatorname*{arg\,min}_{z_{*j}}\Bigl{(}g_{j}(z_{*j})+\frac{\rho}{2}\big{\|}D_{j}z_{*j}-d_{j}+\beta_{j}^{(k)}\big{\|}_{2}^{2}\\ &+\frac{\rho}{2}\big{\|}x^{(k+1)}_{*j}-z_{*j}+\lambda_{*j}^{(k)}\big{\|}_{2}^{2}\Bigr{)}.\end{split}$
+Acknowledgments
+We thank our anonymous reviewers and shepherd for their insightful comments.
+We also thank Yang Zhou, Justin Chiu, Alexander Rush, and Victor Bahl for their valuable feedback and support.
+We extend special thanks to Suvan Chatakondu for enhancing our code after the paper’s acceptance.
+This work is supported in part by ACE, one of the seven centers in JUMP 2.0, a Semiconductor Research Corporation (SRC) program sponsored by DARPA.
+This work is also supported by National Science Foundation grant CCF-2326605 (Expedition in Computing).
+Appendix
+Appendix A Evaluation Setup of Cluster Scheduling
+We detail the experimental setup for the large-scale cluster scheduling evaluation described in §7.1.1.
+Computing resources.
+We collect 456 distinct GPU/CPU resource types from three benchmark sources [26, 36, 19].
+These resource types differ along multiple dimensions such as vendor, interconnect, memory configuration, deployment platform, etc.
+While it is unlikely today for any single cluster to host all 456 types, recent trends in distributed ML workloads [66, 62] suggest the possibility of scheduling ML jobs across global-scale clusters in the future.
+For each resource type, the number of available instances is randomly drawn from the set $\{8,16,24,\ldots,64\}$, reflecting common modern hardware configurations where nodes are provisioned with GPU/CPUs in multiples of eight.
+This setup results in a total of 16,520 GPU/CPU instances in our simulated environment.
+Jobs.
+We generate 2,588 types of ML jobs based on the Notable AI Models dataset [18].
+For each ML model, we synthesize multiple job types encompassing both training and inference tasks under various numerical precisions.
+The number of compute instances requested by each job is sampled from $\{1,2,4,8,16,32\}$, in line with common practices in data and model parallelism.
+Job throughput values are either taken directly from ML benchmarks [26, 36, 19] or, when unavailable, estimated based on each job’s FLOP requirements and the computational capacity of the respective hardware.
+Scheduling Simulator.
+Our evaluation employs Gavel’s ML scheduling simulator [45], with several adjustments.
+Jobs are time-sliced across available resource types but are not co-located, due to the increased size and resource demands of modern ML workloads.
+To ensure that jobs do not grow beyond cluster capacity, we set the average job inter-arrival time to 100 seconds in the Poisson process.
+We begin the simulation from a steady state, maintaining approximately 1,500 to 2,000 active jobs at any given time.
+Job allocation decisions are made every 6 minutes, consistent with Gavel’s default settings, and the simulation runs for 200 scheduling rounds, or 20 hours.